### v5

Abstract
This work aims to improve the sample efficiency of parallel large-scale ranking and selection (R&S) problems by leveraging correlation information. We modify the commonly used "divide and conquer" framework in parallel computing by adding a correlation-based clustering step, transforming it into "clustering and conquer". Analytical results under a symmetric benchmark scenario show that this seemingly simple modification yields an $\mathcal{O}(p)$ reduction in sample complexity for a widely used class of sample-optimal R&S procedures. Our approach enjoys two key advantages: 1) it does not require highly accurate correlation estimation or precise clustering, and 2) it allows for seamless integration with various existing R&S procedures, while achieving optimal sample complexity. Theoretically, we develop a novel gradient analysis framework to analyze sample efficiency and guide the design of large-scale R&S procedures. We also introduce a new parallel clustering algorithm tailored for large-scale scenarios. Finally, in large-scale AI applications such as neural architecture search, our methods demonstrate superior performance.
keywords:
[…]
1 Introduction
Ranking and selection (R&S) aims to identify the best design from a finite set of alternatives, through conducting simulation and learning about their performances (Bechhofer, 1954). It is typically framed into two main formulations: fixed-precision and fixed-budget. Fixed-precision R&S procedures terminate the simulation once a pre-specified level of precision is achieved. Notable examples include the stage-wise algorithm of Rinott (1978) and the KN family (Kim and Nelson, 2001; Hong and Nelson, 2005; Jeff Hong, 2006), among others. In contrast, fixed-budget R&S procedures stop the simulation once a predetermined total simulation budget is exhausted, with the goal of optimizing precision. This category includes algorithms such as OCBA (Chen et al., 2000; Fu et al., 2007), large-deviation-based approaches (Glynn and Juneja, 2004), and algorithms proposed by Chick and Inoue (2001) and Frazier et al. (2008). For a more comprehensive review of the R&S literature, please refer to Chen et al. (2015) and Hong et al. (2021).
In recent years, R&S algorithms have found significant applications in AI-related problems, such as reinforcement learning (Zhu et al., 2024) and Monte Carlo tree search (Liu et al., 2024). This growing body of work has led to a shift in research focus toward large-scale R&S problems, especially in parallel computing environments (Hunter and Nelson, 2017). Here, "large-scale” refers to a large number of alternatives, denoted by $p$. The literature on large-scale R&S can generally be divided into two branches. One branch focuses on addressing the challenges related to parallel computing implementation, such as information communication, synchronization, and workload balancing. Notable works in this branch include Luo et al. (2015) and Ni et al. (2017). The other branch focuses on improving sample efficiency. Recent efforts in this branch primarily aim to modify the inefficient all-pairwise comparison paradigm of classic fully-sequential R&S algorithms such as the KN family. For example, Zhong and Hong (2022) introduce a Knockout Tournament (KT) paradigm (and its fixed-budget version, FBKT (Hong et al., 2022)), which restricts comparisons to "matches" involving only two alternatives at a time. The PASS paradigm of Pei et al. (2022) compares each alternative against a common standard to avoid exhaustive pairwise comparisons.
However, existing R&S algorithms typically assume independence across alternatives, thereby discarding valuable shared information (Eckman and Henderson, 2022). Leveraging such shared information offers a promising direction for improving sample efficiency, yet this remains largely unexplored in the large-scale R&S literature. Existing efforts in utilizing shared information can be broadly categorized into two streams. The first involves using additional contextual or covariate information (L. Salemi et al., 2019; Shen et al., 2021; Du et al., 2024). However, this approach relies on specific problem structures and is not applicable for general-purpose use. The second stream focuses on exploiting correlation or similarity information (Fu et al., 2007; Frazier et al., 2009; Qu et al., 2015; Zhou et al., 2023). However, none of these algorithms is fully suitable for large-scale problems, as they typically require precise estimation of correlation or similarity parameters.
Before introducing our algorithm, we first provide some necessary background: most prominent parallel R&S procedures adopt the divide and conquer framework (Ni et al., 2017; Zhong and Hong, 2022), where alternatives are randomly distributed across processors, then the local best is selected from each processor, and the global best is selected from these local bests. In our work, we extend the traditional divide and conquer framework by adding a correlation-based clustering step. This modified approach, termed the Parallel Correlation Clustering and Conquer (P3C) procedure, clusters alternatives based on their correlation and assigns alternatives from the same cluster to a single processor, rather than distributing them randomly. Both theoretical and empirical evidence show that this simple modification significantly improves sample efficiency. Moreover, P3C enjoys the advantage of not requiring precise estimation of the correlation parameter, as it is sufficient to merely identify which alternatives are highly correlated. Other notable alternative assignment strategies for divide and conquer framework include the seeding approaches (Hong et al., 2022; Li et al., 2024) and the referencing approach Zhong et al. (2025). Moreover, we note that this work studies a general setting without distinguishing the specific sources of correlation. The correlation among alternatives may arise from common random numbers (CRN) introduced in simulation experiments (Chen et al., 2012), or from other forms of correlated sampling.
The intuition behind P3C is that, clustering highly correlated alternatives together can effectively cancel out stochastic fluctuations in the same direction. To be more specific, the reason why R&S requires a large number of samples is due to the random simulation outputs, which occasionally leads to undesirable situations where "good" alternatives may unexpectedly perform worse than "bad" ones. If highly correlated alternatives are grouped together for comparison, their fluctuations tend to align in the same direction to some extent. When a "good" alternative occasionally performs poorly, the "bad" alternatives are likely to show similar declines; conversely, when a "bad" alternative performs better than expected, the "good" alternatives tend to exceed expectations as well. This way, the true ranking is preserved, enabling us to identify the true best alternative with fewer simulations. This concept is similar to the CRN technique, which introduces positive correlation artificially to reduce variance and expedite pairwise comparisons. However, it is important to note that our theoretical analysis is fundamentally different from CRN, as we focus on the global impact of correlation information: for instance, one interesting conclusion from our analysis is that increasing the correlation between a pair of alternatives can increase the probability of selecting other "good" alternatives that are not directly related to this pair.
To formalize the intuition of P3C, we develop a novel gradient analysis framework. In the literature, the probability of correct selection (PCS) is commonly used as a measure of the precision of R&S procedures. It is defined as the probability that the sample average of the "true best" alternative is higher than that of the others. We generalize the classical PCS by replacing the "true best" with any alternative $\tau$, termed as the individual PCS($\tau$), which serves as a probabilistic criterion for assessing the performance of alternative $\tau$. We then analyze the derivative of individual PCS with respect to correlation information to explore a novel "mean-covariance" interaction, as opposed to the widely discussed "mean-variance” tradeoff. The analysis shows that, increasing correlation between alternatives induces an interesting "separation" effect, which probabilistically amplifies good alternatives while suppressing the bad ones. These theoretical insights explain why P3C can enhance sample efficiency thorough correlation-based clustering. Furthermore, by performing gradient analysis with respect to the sample size, we quantify the reduction in sample complexity that correlation-based clustering can bring in parallel computing environments. The gradient analysis established in this paper provides a fundamental framework for studying the sample efficiency and guiding the design of large-scale R&S procedures. Similarly, a related work by Peng et al. (2017) also explores gradient analysis in R&S, but it investigates the impact of "induced correlation", which is induced by variance under the independent assumption, rather than the actual correlation existing between alternatives. Additionally, other works exploring gradient analysis in R&S include Peng et al. (2015) and Zhang et al. (2023).
In recent literature, analyzing the asymptotic behavior of the required total sample size as $p\to\infty$ has become a critical approach for assessing the sample efficiency of large-scale R&S procedures (Zhong and Hong, 2022; Hong et al., 2022; Li et al., 2024). As $p\to\infty$, the theoretical lowest growth rate of the required total sample size to achieve non-zero precision asymptotically is $\mathcal{O}(p)$, and a R&S procedure that achieves this $\mathcal{O}(p)$ sample complexity is referred to as sample-optimal. Known sample-optimal procedures include the median elimination (ME) procedure (Even-Dar et al., 2006), the KT procedure and its fixed-budget version FBKT, as well as the greedy procedure in Li et al. (2024). As $p$ grows large, the proposed P3C performs multiple rounds of "clustering and conquer,” resembling the knockout tournament scheme, and can also achieve sample optimality when combined with classic fixed-budget and fixed-precision R&S algorithms. Moreover, under a symmetric benchmark scenario, the reduction in sample complexity achieved by P3C is $\mathcal{O}(p)$. This implies that, although sample-optimal R&S procedures already achieve the lowest $\mathcal{O}(p)$ growth rate in complexity, P3C can further reduce the slope.
The main contributions of this paper are as follows. In Section 3, we propose the P3C procedure and in Section 4, we develop a novel gradient-based framework to analyze and guide sample-efficient R&S design. This framework reveals a key mean-covariance interaction and shows that correlation-based clustering in P3C enables $\mathcal{O}(p)$ sample complexity reduction under a symmetric benchmark scenario. In Section 5, we introduce a parallelizable few-shot clustering algorithm.
2 Problem Formulation
Let $\mathcal{P}=\{1,2,\ldots,p\}$ denote the index set for all $p$ alternatives. We adopt a frequentist framework, and the output of alternative $i\in\mathcal{P}$ is a random variable $X_{i}$. We assume that the population distribution of the random vector $(X_{1},X_{2},\ldots,X_{p})$ is multivariate normal $N(\bm{\mu},\Sigma_{p\times p})$, where $\bm{\mu}=(\mu_{1},\cdots,\mu_{p})$ is the mean vector and $\Sigma_{p\times p}$ is the covariance matrix. Let $x_{ij}$ denote the $j$th simulation observation of alternative $i$. The observation vectors $(x_{1j},x_{2j},\cdots,x_{pj})\sim N(\bm{\mu},\Sigma_{p\times p})$ are independently and identically distributed. We assume that the covariance between $x_{im}$ and $x_{jn}$ is given by $\mathrm{cov}(x_{im},x_{jn})=0$ if $m\neq n$ and $\mathrm{cov}(x_{im},x_{jn})=\mathrm{cov}(X_{i},X_{j})$ if $m=n$, where $\mathrm{cov}(X,Y)$ denotes the covariance between the random variables $X$ and $Y$. Let $N_{i}$ denote the total sample size allocated to alternative $i$ and ${\bar{x}}_{i}$ be the sample average. Then $\mathrm{cov}\left({\bar{x}}_{i},{\bar{x}}_{j}\right)=\frac{\mathrm{cov}\left(X_{i},X_{j}\right)}{N_{i,j}}$, where $N_{i,j}\triangleq\max(N_{i},N_{j})$. The correlation among alternatives may arise from any form of correlated sampling. In practice, some alternatives are more correlated than others, leading to a natural clustering structure in which within-cluster correlations are stronger than between-cluster correlations. The correlation structure, as well as the underlying cluster partition, is unknown and needs to be estimated using observed data.
Let $[m]$ denote the index of the alternative with $m$-largest mean, i.e., $\mu_{[1]}>\mu_{[2]}>\cdots>\mu_{[p]}$. The objective of R&S is to identify the true best alternative $\left[1\right]=\mathop{\arg\max}\limits_{i\in\mathcal{P}}\ \mu_{i}$, under either a fixed budget constraint or a fixed precision constraint. Since the selection policy is typically specified as selecting $\arg\max_{i\in\mathcal{P}}{\bar{x}}_{i}$, the corresponding precision metric is the traditional PCS, defined as $\text{PCS}_{\text{trad}}\triangleq P\left({\bar{x}}_{[1]}>{\bar{x}}_{j},j\neq[1]\right)$. In this work, to facilitate a more general theory, for any alternative $\tau\in\mathcal{P}$, we define the individual PCS of $\tau$ as PCS(τ)≜P(¯x_τ>¯x_j,j≠τ), which naturally generalizes the traditional PCS as $\text{PCS}_{\text{trad}}=\text{PCS}\left([1]\right)$. According to Hong et al. (2021), the statistical meaning of $\text{PCS}\left(\tau\right)$ is the probability of rejecting the null hypothesis that $\tau$ is not the true best", making it a useful metric for evaluating the performance of $\tau$.
Recently, analyzing the asymptotic behavior as $p\to\infty$ has emerged as a critical approach for assessing the theoretical performance of large-scale R&S procedures (Zhong and Hong, 2022; Hong et al., 2022; Li et al., 2024). As $p\to\infty$, the theoretical lowest growth rate of the required total sample size to deliver a non-zero precision asymptotically is $\mathcal{O}(p)$. A R&S procedure that achieves this $\mathcal{O}(p)$ sample complexity is referred to as sample-optimal in this paper. When examining the asymptotic behavior as $p\to\infty$, we assume the same regime as Li et al. (2024): the index of the true best $[1]$ remains unchanged, and the difference between $[1]$ and $[2]$ remains above a positive constant and the covariance between any two alternatives is upper bounded by a constant.
3 The Framework of P3C: From "Divide and Conquer" to "Clustering and Conquer"
"Divide and conquer" has long served as the foundational framework for many mainstream parallel R&S procedures. In this framework, alternatives are randomly distributed across different processors, where each processor selects its local best alternative, and the global best is then selected from these local bests. We extend this framework by introducing a "clustering and conquer" strategy, formally termed Parallel Correlation Clustering and Conquer (P3C). The key distinction in P3C is an additional step of correlation-based clustering, which groups highly correlated alternatives and assigns each cluster to a single processor, as opposed to random assignment. This simple modification leads to a significant improvement in sample efficiency.
Specifically, as outlined in Algorithm 1 and Figure 1, P3C starts with an initialization Stage 0. Then, in Stage 1, we continue sampling and then cluster the alternatives based on estimated correlations. The sample size can be predetermined based on the required clustering accuracy. Additionally, the novel few-shot clustering algorithm $\mathcal{A}\mathcal{C}^{+}$, detailed in Section 5, enables efficient parallelization of both Stage 0 and Stage 1 in large-scale problems. Upon completion of Stage 1, the alternatives of the same cluster are sent to a single processor. During Stage 2, we perform R&S to select the local best within each cluster. Here, the choice of R&S algorithm is flexible and depends on the specific problem formulation. For example, the KT or KN family can be applied in the fixed-precision P3C (denoted as P3C-KT and P3C-KN, respectively). In Stage 3, we continue R&S within these remaining local bests to select the final winner.
Figure 1: Parallel correlation clustering and conquer.
However, when $p$ is large, the number of remaining local bests after Stage 2 may still be substantial. In such cases, P3C employs repeated rounds of "clustering and conquer" (i.e., Stages 1 and 2) until the number of remaining alternatives becomes manageable for a single processor. In practice, the maximum number of alternatives that a single processor can handle, denoted by $p_{m}\geq 2$, is pre-estimated based on computational capacity. In the fixed-precision R&S, let the final desired precision be $1-\alpha$, where $\alpha$ is the false elimination probability (FEP). The FEP $\alpha_{2}^{r}$ for each processor in the $r$-th round of "clustering and conquer", and the FEP $\alpha_{3}$ for Stage 3 in P3C must satisfy $\sum_{r}\alpha_{2}^{r}+\alpha_{3}=\alpha$. We set $\alpha_{2}^{r}=\frac{\alpha-\alpha_{3}}{2^{r}}$. This multi-round structure closely resembles the knockout tournament paradigm, and consequently, as shown in the following Proposition 1, P3C-KN and P3C-KT can also achieve sample optimality as $p\to\infty$, provided cluster sizes are bounded by a finite $p_{m}$. In the fixed-budget R&S, P3C also adopts a multi-round structure when $p$ is large and is shown to achieve sample optimality. The main body of this paper focuses on the fixed-precision setting, so we defer the detailed discussion of fixed-budget P3C to the supplementary materials.
Proposition 1 (Sample Optimality of P3C-KN and P3C-KT)
Let $N_{\mathrm{P3C-KN}}$ and $N_{\mathrm{P3C-KT}}$ denote the required total sample sizes to achieve $\mathrm{PCS_{trad}}\geq 1-\alpha$ using $\mathrm{P3C-KN}$ and $\mathrm{P3C-KT}$, respectively. If the number of alternatives within each cluster is upper bounded by a constant $p_{m}<\infty$, then $\mathbb{E}(N_{\mathrm{P3C-KN}})=\mathcal{O}(p)$ and $\mathbb{E}(N_{\mathrm{P3C-KT}})=\mathcal{O}(p)$ as $p\to\infty$.
Note that this paper focuses on analyzing the sample efficiency of the P3C, while many challenges in parallel implementation, such as master-worker task coordination, synchronization, and workload balancing, are beyond our scope. For comprehensive solutions to these issues, readers are encouraged to refer to the prior works of Luo et al. (2015) and Ni et al. (2017). It is also important to note that, in our setting, no additional structural assumptions are imposed, and highly correlated alternatives do not necessarily have similar means. Therefore, grouping highly correlated alternatives does not exacerbate potential drawbacks associated with clustering mean-similar alternatives. In addition, the clustering algorithm $\mathcal{A}\mathcal{C}^{+}$ described in Section 5 incorporates specific steps to further prevent such cases.
Algorithm 1 Parallel Correlation Clustering and Conquer (P3C)
Stage 0 (Initialization): Simulate each alternative at least $N_{0}\geq 3$ times and estimate the covariance matrix $\Sigma$.
Stage 1 (Clustering): Continue sampling and apply the correlation-based clustering algorithm $\mathcal{A}\mathcal{C}^{+}$ to partition all alternatives into several clusters. Assign all alternatives within the same cluster to a single processor.
Stage 2 (Conquer): Perform R&S within each cluster to select the local best.
Stage 3 (Final Comparison): If the number of remaining alternatives in contention satisfies $\leq p_{m}$, perform R&S to select the global optimal alternative. Otherwise, repeat Stage 1 and Stage 2 for the remaining alternatives.
4 Theoretical Analysis: Understanding Mean-Correlation Interactions
In this section, we present a comprehensive analysis of P3C from a novel gradient-based perspective. Viewing $\text{PCS}\left(\tau;\bm{\mu},\Sigma,\{N_{i}\}_{i\in\mathcal{P}}\right)$ as a function of mean, covariance and the allocation of sample sizes, we analyze its derivatives to gain insights for designing large-scale R&S procedures. Section 4.1 examines the interaction between mean and correlation, revealing an interesting "separation" effect that explains why correlation-based clustering in P3C enhances sample efficiency. Section 4.2 quantifies the sample complexity reduction brought about by P3C. Our theoretical analysis based on individual $\text{PCS}(\tau)$ is more general and encompasses $\text{PCS}_{\text{trad}}$ by setting $\tau=[1]$. Note that the total sample sizes $\{N_{i}\}_{i\in\mathcal{P}}$ are random variables that depend on the sampling policy and past observations. In our gradient-based analysis, we temporarily ignore the stochastic nature of $\{N_{i}\}_{i\in\mathcal{P}}$, as the goal is to understand how increases or decreases in sample sizes interact with the correlation structure to provide insights into correlation-based clustering, rather than to design a dynamic sample allocation algorithm.
Before proceeding, we introduce additional notations. Let $\sigma^{2}_{i}$ denote the (unknown) variance of $X_{i}$ and $r_{ij}$ denote the (unknown) Pearson correlation coefficient between $X_{i}$ and $X_{j}$. For any alternative $\tau\in\mathcal{P}$, $\text{PCS}\left(\tau\right)$ can be rewritten as $P(y^{\tau}_{1}>-d^{\tau}_{1},...,y^{\tau}_{p}>-d^{\tau}_{p})$, where $y^{\tau}_{i}=\frac{\bar{x}_{\tau}-{\bar{x}}_{i}-(\mu_{\tau}-\mu_{i})}{\sqrt{\lambda^{\tau}_{i}}}$, $d^{\tau}_{i}=\frac{\mu_{\tau}-\mu_{i}}{\sqrt{\lambda^{\tau}_{i}}}$, $\lambda^{\tau}_{i}=\mathrm{var}({\bar{x}}_{\tau}-{\bar{x}}_{i})=\frac{\sigma^{2}_{\tau}}{N_{\tau}}+\frac{\sigma^{2}_{i}}{N_{i}}-2\frac{\mathrm{cov}(X_{\tau},X_{i})}{N_{\tau,i}}$, $i\in\mathcal{P}\setminus\{\tau\}$. The vector $\bm{y^{\tau}}=(y^{\tau}_{1},...,y^{\tau}_{\tau-1},y^{\tau}_{\tau+1},...,y^{\tau}_{p})$ follows distribution $N(0,\Phi^{\tau})$, with the covariance matrix $\Phi^{\tau}=\big(\tilde{r}^{\tau}_{i,j}\big)_{(p-1)\times(p-1)}$ where $i,j\in\mathcal{P}\setminus\{\tau\}$. The diagonal elements of $\Phi^{\tau}$ are 1 and $|\tilde{r}^{\tau}_{i,j}|\leq 1$. We further note that the theoretical results are highly technical and involve intricate interactions between sample allocation and distribution parameters. To provide more concise results and clearer insights, we impose only a mild assumption that $\tilde{r}^{\tau}_{i,j}$ is bounded.
Assumption 1 (Moderate Correlation)
$\forall i,j\in\mathcal{P}\setminus\{\tau\}$ such that $d_{i}^{\tau}\neq 0$, the following conditions hold: $\mathrm{(a)}$ $|\tilde{r}^{\tau}_{i,j}|<\big|\frac{d^{\tau}_{j}}{d^{\tau}_{i}}\big|$; $\mathrm{(b)}$ $\tilde{r}^{\tau}_{i,j}<\big|\frac{d^{\tau}_{j}}{d^{\tau}_{i}}\big|^{2}$.
Different theorems may incorporate one or more of the above assumptions as needed. In the supplementary material, we show that these assumptions are readily satisfied for prominent R&S algorithms, as long as the correlations between alternatives are not extremely large (always hold under independence).
4.1 Interaction and Impact of Mean and Correlation on PCS
We first present the technical results in Subsection 4.1.1, and then, in Subsection 4.1.2, we intuitively explain the underlying "separation" effect and show how P3C leverages this insight to accelerate R&S process. The proof of the technical results is provided in Appendix A.
4.1.1 The derivative of PCS with respect to mean and correlation information.
To begin with, the following fundamental property establishes that $\mathrm{PCS}(\tau)$ prioritizes mean information and increases monotonically with respect to the mean $\mu_{\tau}$ under any correlation configuration.
Theorem 1
$\mathrm{PCS}\left(\tau\right)$ is differentiable with respect to $\bm{\mu}=(\mu_{1},...,\mu_{p})$ and $\frac{\partial\mathrm{PCS}\left(\tau\right)}{{\partial\mu}_{\tau}}>0$.
Then we examine a more refined interaction between mean and correlation by analyzing the gradient with respect to correlation.
Theorem 2
Let $\tau=[m]$, we define $\mathcal{I}^{+}(\tau)\triangleq\{[i]:i<m\}$ and $\mathcal{I}^{-}(\tau)\triangleq\{[i]:i>m\}$, which are the index sets of the alternatives with larger and smaller mean than $\tau$, respectively. We denote the sign of a real number as $\mathrm{sign}(\cdot)$. Then $\forall i\in\mathcal{P}\setminus\{\tau\}$,
$\frac{\partial PCS\left(\tau\right)}{\partial r_{\tau i}}=D_{i}^{\tau}+I_{i}^{\tau},$
$\text{where}\ D_{i}^{\tau}\triangleq\frac{\partial\mathrm{PCS(\tau)}}{\partial d_{i}^{\tau}}\frac{\partial d_{i}{}^{\tau}}{\partial r_{\tau i}},\ \ I_{i}^{\tau}=\sum_{j\in\{1,\cdots,p\}\setminus\{i,\tau\}}I_{i}^{\tau,j}\triangleq\sum_{j\in\{1,\cdots,p\}\setminus\{i,\tau\}}2\frac{\partial\text{PCS}(\tau)}{\partial{{\widetilde{r}}_{i,j}}^{\tau}}\frac{\partial\widetilde{r}_{i,j}^{\tau}}{\partial r_{\tau i}}.$
$\mathrm{(a)}$ $D_{i}^{\tau}>0$ for $i\in\mathcal{I}^{-}(\tau)$ and $D_{i}^{\tau}<0$ for $i\in\mathcal{I}^{+}(\tau)$;
$\mathrm{(b)}$ $\mathrm{sign}(I_{i}^{\tau,j})=\mathrm{sign}\left(-\sigma_{i}^{2}(N_{i})^{-1}+{r_{\tau,i}\sigma_{\tau}\sigma_{i}}{(N_{\tau,i})^{-1}}-{r_{\tau,j}\sigma_{\tau}\sigma_{j}}{(N_{\tau,j})^{-1}}+{r_{i,j}\sigma_{i}\sigma_{j}}{(N_{i,j})^{-1}}\right).$
Remark 1
The specific expressions of terms $D_{i}^{\tau}$ and $I_{i}^{\tau}$ can be found in the appendix. The analysis presented in this paper does not specify the computation of these terms but can help gain insights through the magnitude and sign of each term. To complete the analysis of the "mean-covariance" interaction, we also present the impact of variance information in the appendix.
As shown in Theorem 2, the derivative $\frac{\partial PCS\left(\tau\right)}{\partial r_{i\tau}}$ is composed of two parts: the mean-determined (MD) term $D_{i}^{\tau}$ and the mean-independent (MI) term $I_{i}^{\tau}$. The sign of the $D_{i}^{\tau}$ depends on whether the mean $\mu_{i}$ is larger or smaller than $\mu_{\tau}$, whereas the sign of $I_{i}^{\tau}$ is independent of the mean. Determining the sign of $I_{i}^{\tau}$ is generally complex and unclear. Therefore, it is necessary to establish the magnitudes of $D_{i}^{\tau}$ and $I_{i}^{\tau}$ terms to ascertain which term dominates. For a given vector $\bm{d}$ and an index set $S$, we write $\bm{d}_{S}$ the vector composed of the components from $\bm{d}$ with indices $S$. Similar notations $\Sigma_{S}$ are used for submatrices of matrix $\Sigma$. Let $N$ denote the total sample size, i.e., $N=\sum_{i\in\mathcal{P}}N_{i}$.
Assumption 2
$N_{i}=\mathcal{O}\left(N_{j}\right)$ as $N\to\infty$, $\forall i,j\in\mathcal{P}$.
Remark 2
In Assumption 2 and the following Corollary 1, we consider the asymptotic regime where $N\to\infty$ while $p$ is fixed. For sampling policies that may eliminate certain alternatives, an additional $\epsilon$-greedy step can be incorporated to ensure that Assumption 2 holds; that is, with probability $\epsilon$, an alternative is selected uniformly from $\mathcal{P}$ and sampled; otherwise, the original sampling policy is followed. The $\epsilon$ can be chosen arbitrarily small so that its practical impact on the overall allocation is negligible.
Corollary 1
$\forall i\in\mathcal{P}\setminus\{\tau\}$, we define $\mathcal{S}_{-i}^{+}(\tau)\triangleq\{j\in\mathcal{P}\setminus\{i,\tau\}|-\widetilde{d}_{j}^{\tau}>0\}$, $\widetilde{d}_{j}^{\tau}\triangleq-d_{j}^{\tau}+d_{i}^{\tau}\widetilde{r}^{\tau}_{i,j}$, $j\in\mathcal{P}\setminus\{\tau,i\}$. If Assumption 2 holds, as $N\to\infty$, then
$\mathrm{(a)}$
$\left|D_{i}^{\tau}\right|=\mathcal{O}(e^{-\mathscr{D}^{\tau}_{i}N})\triangleq\mathcal{O}\bigg(\exp\big(-\frac{(d_{i}^{\tau})^{2}+Q^{\tau}_{S}}{2}\big)\bigg),$ (1)
$\left|I_{i}^{\tau}\right|=\mathcal{O}(e^{-\mathscr{I}^{\tau}_{i}N})\triangleq\mathcal{O}\bigg(\exp{\big(-\min_{j\neq i}\frac{(d_{i}^{\tau})^{2}+(d_{j}^{\tau})^{2}}{2(1+|{{\widetilde{r}}_{i,j}}^{\tau}|)}\big)}\bigg),$ (2)
where $Q^{\tau}_{S}\triangleq\min_{\bm{x}\geq\bm{d}_{\mathcal{S}_{-i}^{+}(\tau)}}\langle\bm{x},(\Sigma^{Z}_{\mathcal{S}_{-i}^{+}(\tau)})^{-1}\bm{x}\rangle$, $\bm{x}=(x_{1},\cdots,x_{|\mathcal{S}_{-i}^{+}(\tau)|})\in\mathbb{R}^{|\mathcal{S}_{-i}^{+}(\tau)|}$, $\bm{d}=(-\widetilde{d}_{1}^{\tau},\cdots,-\widetilde{d}_{p}^{\tau})\in\mathbb{R}^{p-2}$ and $\Sigma^{Z}$ is a covariance matrix of $p$ transformed variables which are given in the appendix;
[…]
Some key observations follow. First, both the MD and MI terms will decrease exponentially to 0 as the total sample size $N$ increases. Whether the MD terms or MI terms dominate depends on the relative magnitudes of decay rates $\mathscr{D}^{\tau}_{i}$ and $\mathscr{I}^{\tau}_{i}$. Second, combining Corollary 1(b) and Corollary 1(c), let $\tau=[m]$. When $m=1$, i.e., when $\text{PCS}(\tau)=\text{PCS}_{\text{trad}}$, the MD term is dominant. As the ranking $m$ increases, the MI term gradually surpasses the MD term. While the exact ranking $m$ at which the MI term becomes the dominant component is unclear, we can at least confirm that there exists a subset $\mathcal{P}^{\prime}\subseteq\mathcal{P}$ containing $[1]$, within which the MD term remains dominant for all alternatives. The alternatives in $\mathcal{P}^{\prime}$ are top-ranked "good" alternatives, and in practice, users are typically only interested in these "good" ones. Therefore, in the intuitive analysis in Section 4.1.2, we assume that the MD term dominates for any alternative under consideration (this assumption is used exclusively in the intuitive analysis below).
4.1.2 "Separation" effect
Figure 2: An illustrative example with 7 alternatives, where the top 3 alternatives $[1]$, $[2]$ and $[3]$ exhibit high mean performances, and $[5]$, $[6]$ and $[7]$ exhibit low mean performances.
Based on the aforementioned technical results, we are now ready to show the underlying insight. According to the Theorem 2(a), focusing solely on the MD term, the correlations between $\tau$ and alternatives in $\mathcal{I}^{-}(\tau)$ impose a cumulative positive effect on $\text{PCS}\left(\tau\right)$, and conversely, correlations between $\tau$ and those alternatives in $\mathcal{I}^{+}(\tau)$ impose a cumulative negative effect. Consequently, the impact of correlation on $\text{PCS}\left(\tau\right)$ hinges on the ranking of $\tau$. If $\tau$ is near the top, the number of alternatives in $\mathcal{I}^{-}(\tau)$ dominates, and then the cumulative positive effects surpass the cumulative negative effects. Increasing the correlations $\{r_{\tau i}\}_{i\in\mathcal{P}\setminus\{\tau\}}$ is more likely to improve $\text{PCS}(\tau)$ (specifically, when $\tau=[1]$ and $\mathcal{I}^{+}(\tau)=\emptyset$, then increasing the correlation will certainly improve $\text{PCS}_{\text{trad}}$). On the contrary, if the alternative $\tau$’s mean performance is poor, the influence of correlation is reversed, and increasing the correlation around $\tau$ will decrease $\text{PCS}(\tau)$.
In summary, as shown in regions I and IV of Figure 2, increasing the correlation level among the alternatives can induce a "separation" effect, which may "amplify" good alternatives and "suppress" bad alternatives in terms of individual PCS. This makes the distinction between good and bad alternatives more pronounced, thereby accelerating the R&S procedure. Conversely, decreasing the overall correlation leads to an undesirable aggregation phenomenon (regions II and III). In practice, we cannot directly modify the correlation parameters within the true distribution of alternatives. However, in a parallel computing environment, we can enhance the local correlation level in each processor using clustering techniques. This is the rationale behind P3C’s use of correlation-based clustering to enhance sample efficiency.
Before concluding this section, we present the following lemma, which will be used in the next section as a supplement to Theorem 2.
Lemma 1
$\forall i,j\in\mathcal{P}\setminus\{\tau\}$, $\frac{\partial\text{PCS}\left(\tau\right)}{\partial r_{ij}}\geq 0$. If $\mathscr{D}^{\tau}_{i}<\mathscr{I}^{\tau}_{i}$, then $\frac{\partial\text{PCS}\left(\tau\right)}{\partial r_{ij}}=o\bigg(|\frac{\partial\text{PCS}\left(\tau\right)}{\partial r_{\tau i}}|\bigg)$.
This lemma characterizes the impact of correlations $\{r_{ij}\}_{i,j\in\mathcal{P}\setminus\{\tau\}}$, which are not directly associated with $\tau$. These correlations in the surrounding environment always have a non-negative impact on $\text{PCS}(\tau)$, regardless of mean information. This further complements the argument that clustering correlated alternatives can accelerate the R&S procedure. Additionally, if $\tau$ is a "good" alternative within $\mathcal{P}^{\prime}$, meaning that $\mathscr{D}^{\tau}_{i}<\mathscr{I}^{\tau}_{i}$, the impact of $\{r_{ij}\}_{i,j\in\mathcal{P}\setminus\{\tau\}}$ is negligible compared to $\{r_{\tau i}\}_{i\in\mathcal{P}\setminus\{\tau\}}$, so it will not affect the aforementioned "separation" phenomenon induced by $\{r_{\tau i}\}_{i\in\mathcal{P}\setminus\{\tau\}}$.
4.2 Quantifying the Sample Complexity Reduction
After understanding how P3C utilizes correlation-based clustering to accelerate the R&S process, we now proceed to quantify the reduced sample complexity in fixed-precision R&S. To gain analytical clarity, we focus on the symmetric benchmark scenario, a simplified yet representative setting that enables a clean derivation highlighting the order of sample complexity reduction, without being obscured by intricate expressions or case-specific constants.
In this scenario, we assume that the $p$ alternatives come from $k$ non-overlapping clusters based on correlation: $\mathcal{G}_{1},\ldots,\mathcal{G}_{k}$, with cardinality $|\mathcal{G}_{j}|=p_{j}$. $G$ is a mapping from $\mathcal{P}$ to $\{1,2,\ldots,k\}$, where $G\left(i\right)=j$ if alternative $i$ belongs to $\mathcal{G}_{j}$. $\Pi=(G\left(1\right),G\left(2\right),\ldots,G(p))\in\mathbb{R}^{p}$ represents the true cluster partition. The independent case is also included by setting $p=k$. Alternatives within the same cluster are more highly correlated than those across clusters, i.e., the correlation structure satisfies the following Assumption 3.
Assumption 3
$r_{ab}>r_{cd}$ for all $a,b,c,d\in\mathcal{P}$ such that $G(a)=G(b)$ and $G(c)\neq G(d)$.
We consider a parallel computing environment with $k\leq p$ available processors and compare two distinct strategies: ($\mathcal{R}$) randomly assigning an equal number of alternatives to each processor, as in "divide and conquer" procedures; and ($\mathcal{C}$) adopting a correlation-based clustering approach, assigning each cluster to a single processor, as in P3C. In the symmetric benchmark regime, we consider a stylized case in which each cluster $\mathcal{G}_{j}$ ($j=1,\cdots,k$) contains one local best alternative $\tau_{j}$, which has the highest mean, along with $\frac{p}{k}$ suboptimal alternatives. The true clustering label is unknown, but we assume knowledge of the indices of the local bests $\{\tau_{j}\}_{j=1,\cdots,k}$. Correlation coefficients within the same cluster are $R$, while those between different clusters are $r$, with a difference of $\Delta r=R-r$. Each local best is placed on a dedicated processor and remains fixed, without being relocated to another processor in either strategy $\mathcal{R}$ or $\mathcal{C}$. All $k$ local bests share the same mean and variance. The remaining $p$ suboptimal alternatives, which also have identical means and variances, are distributed across processors according to strategy $\mathcal{R}$ or $\mathcal{C}$. Then, samples are simulated until $\text{PCS}_{\text{trad}}>1-\alpha$ is achieved. The required total sample sizes in strategies $\mathcal{R}$ and $\mathcal{C}$ are denoted as $N_{R}$ and $N_{C}$, respectively.
Due to the limited number of samples used for learning correlation information, the clustering accuracy in strategy $\mathcal{C}$ may be less than $1$. Given the randomness of simulation outputs, we introduce probability of correct clustering (PCC) to measure the statistical guarantee of clustering quality, defined as PCC≜P(Π_n=Π), where $\Pi_{n}=\left(G_{n}(1),\ldots,G_{n}(p)\right)\in\mathbb{R}^{p}$ is the partition result obtained by the employed clustering algorithm using $n$ samples. We omit $n$ in the notation of $\mathrm{PCC}$. Then, the following Theorem 3 quantifies the sample complexity reduction caused by correlation-based clustering. The proof is provided in Appendix B.
Assumption 4
$\mathrm{PCS}_{\mathrm{trad}}$ is a monotonically increasing and concave function with respect to the total sample size $N$.
Theorem 3
[…]
$\mathbb{E}(N_{R}-N_{C})\geq\gamma\Delta r\bigg(\mathrm{PCC}-\frac{1}{k}\bigg)p,$ (3)
where $\gamma=\frac{\partial\mathrm{PCS}_{\mathrm{trad}}^{1}\left(\Sigma^{\prime},N_{\xi}\right)}{\partial r_{i_{0},\tau}}\big(\frac{\partial\mathrm{PCS}_{\mathrm{trad}}^{1}\left(\Sigma,N_{\xi}\right)}{\partial N}\big)^{-1}$, $N_{\xi}=N_{0}+\xi(N_{C}-N_{0})$, with $\mathrm{PCS}_{\mathrm{trad}}^{1}$ being the local $\mathrm{PCS}_{\mathrm{trad}}$ of the first processor, $\Sigma$ and $\Sigma^{\prime}$ being two covariance matrices, $i_{0}\in\mathcal{G}_{1}\setminus\{\tau_{1}\}$ and $N_{0}$ being the initialization sample size.
Remark 3
Assumption 4 implies that, in the benchmark scenario, a sensible sampling policy is adopted such that $\mathrm{PCS}_{\mathrm{trad}}$ improves as the total sample size increases. The concavity assumption means that as the sample size increases, the marginal improvement in $\mathrm{PCS}_{\mathrm{trad}}$ (upper bounded by 1) gradually diminishes. This assumption can be removed, yielding similar but less concise results. The employed sampling policy is reflected in the denominator of $\gamma$ (i.e., $\frac{\partial\mathrm{PCS}_{\mathrm{trad}}}{\partial N}$), influencing how fast $\mathrm{PCS}_{\mathrm{trad}}$ grows with the total sample size.
Some key observations of Theorem 3 follow. First, correlation-based clustering guarantees a positive reduction in sample complexity as long as $\mathrm{PCC}>\frac{1}{k}$. Second, for the class of sample-optimal R&S procedures which achieve a sample complexity of $\mathcal{O}(p)$ as $p\to\infty$, the reduction attained by correlation-based clustering is also $\mathcal{O}(p)$, since $\gamma$ has been proven to be at least $\mathcal{O}(1)$ for this class (see Appendix B for the proof). This implies that, although sample-optimal R&S procedures already achieve the lowest linear growth rate, P3C can further reduce the slope of the line.
Although the analysis in this subsection is based on a stylized benchmark scenario for clarity, the proof reveals that the $\mathcal{O}(p)$ rate remains valid even without certain assumptions (e.g., equal cluster sizes, uniform correlation difference $\Delta r$). The underlying intuition is that the total improvement essentially accumulates linearly over $p$ alternatives. The intuition of the proof is summarized by the following informal equation, which is established within each processor:
$\Delta\mathrm{PCS}_{\mathrm{trad}}=\underbrace{\frac{\partial\mathrm{PCS}_{\mathrm{trad}}}{\partial N}(N_{R}-N_{0})}_{\text{Strategy}\ \mathcal{R}}=\underbrace{\overbrace{\frac{\partial\mathrm{PCS}_{\mathrm{trad}}}{\partial\Sigma}(\Sigma^{\prime}-\Sigma)}^{\text{Sample Complexity Reduction}}+\frac{\partial\mathrm{PCS}_{\mathrm{trad}}}{\partial N}(N_{C}-N_{0})}_{\text{Strategy}\ \mathcal{C}}.$ (4)
Equation (4) is derived using the mean value theorem, and we omit the exact evaluation points for the derivatives $\frac{\partial\mathrm{PCS}_{\mathrm{trad}}(\cdot)}{\partial N}$ and $\frac{\partial\mathrm{PCS}_{\mathrm{trad}}(\cdot)}{\partial\Sigma}$, though it is worth noting that these points differ between terms. The term $\frac{\partial\mathrm{PCS}_{\mathrm{trad}}}{\partial\Sigma}$ is used here as an intuitive representation of the sensitivity to changes in correlation structure, as described in Theorem 2 and Lemma 1. This equation implies that, both increasing the sample size and adjusting the correlation structure contribute to the improvement in $\mathrm{PCS}_{\mathrm{trad}}$. Strategy $\mathcal{R}$ directly increases the sample size from $N_{0}$ to $N_{R}$, resulting in an improvement of $\Delta\mathrm{PCS}_{\mathrm{trad}}$. In contrast, Strategy $\mathcal{C}$ first modifies the correlation structure from $\Sigma$ to $\Sigma^{\prime}$ through a clustering step, which immediately improves $\mathrm{PCS}_{\mathrm{trad}}$. Therefore, a smaller sample size $N_{C}$ is sufficient to achieve the same precision as Strategy $\mathcal{R}$. The improvement contributed by changes in the correlation structure directly accounts for the reduction in sample complexity, and this improvement accumulates linearly as the number $p$ increases. Notice that, in the case where alternatives are independent, correlation-based clustering brings no improvement.
5 Large-Scale Alternative Clustering
While PCC does not directly determine the final statistical precision of R&S, a higher PCC, as shown in Theorem 3, can facilitate greater sample savings. Nevertheless, clustering in large-scale problems presents challenges, both computationally and statistically. To address these challenges and strive for a high PCC, inspired by a well-known few-shot learning approach in the deep learning literature, namely, Prototypical Networks (Snell et al., 2017), we propose a parallel alternative clustering algorithm $\mathcal{A}\mathcal{C}^{+}$. Following this, we detail the methodology for calculating the PCC.
5.1 Few-shot Alternative Clustering
In P3C, we cluster alternatives instead of their simulation outputs, a concept known as variable clustering in statistical literature. A commonly used algorithm for this purpose is the hierarchical clustering algorithm (Jolliffe, 1972), denoted as $\mathcal{A}\mathcal{C}$ (see pseudocode in the supplementary material). The framework of $\mathcal{A}\mathcal{C}$ is as follows. Suppose the number of clusters $k$ is known and each alternative has a sample size of $n$. Initially, each alternative is treated as an individual group. In each iteration, the two groups with the maximum similarity $R$ are merged until $k$ groups remain. The (empirical) similarity between two groups of alternatives, $G^{1}$ and $G^{2}$, is quantified by $R\left(G^{1},G^{2}\right)=\max\limits_{i\in G^{1},j\in G^{2}}\hat{r}_{ij}^{n}$, where $\hat{r}_{ij}^{n}$ denotes the estimated $r_{ij}$ using $n$ samples. However, $\mathcal{A}\mathcal{C}$ still requires estimating the entire $p\times p$ correlation matrix, which necessitates a large sample size, and more importantly, cannot be parallelized.
The few-shot $\mathcal{A}\mathcal{C}^{+}$ algorithm resolves the above issues. As shown in Algorithm 2, initially, $\mathcal{P}$ is split into two sets: the support set $\mathcal{P}_{s}$ and the query set $\mathcal{P}_{q}$, with sizes $p_{s}$ and $p_{q}$ respectively. $p_{s}$ is set to be greater than $k$ but much smaller than $p$. Next, alternatives in $\mathcal{P}_{s}$ are grouped into $k$ clusters using the $\mathcal{A}\mathcal{C}$ algorithm, which can be efficiently handled on a single processor due to the moderate size of $p_{s}$. In each cluster $\mathcal{G}_{j}$ ($j=1,2,\cdots,k$), one representative alternative $\tau_{j}$ is chosen as the "prototype" of the cluster. Finally, each alternative in the query set $\mathcal{P}_{q}$ is assigned to an existing cluster by identifying the most correlated prototype. This matching process can be parallelized by sending a copy of the $k$ prototypes to each processor and then randomly assigning alternatives in $\mathcal{P}_{q}$ to different processors. Given that $k\ll p$ typically, the additional simulation cost for these copies is negligible. $\mathcal{A}\mathcal{C}^{+}$ eliminates the need for estimating the entire correlation matrix and only requires estimating a small submatrix.
As for the selection of the "prototype" within each cluster $\mathcal{G}_{j}$, the process is as follows. (i) Apply principal component analysis (PCA) to the cluster $\mathcal{G}_{j}$ to identify the first principal component $PC_{\mathcal{G}_{j}}$, which is a synthetic variable given by $PC_{\mathcal{G}_{j}}=\Sigma_{i\in\mathcal{G}_{j}}k_{ij}X_{i}$. Detailed calculations can be found in standard machine learning textbooks. In variable clustering literature, $PC_{\mathcal{G}_{j}}$ is often termed as the "latent component" of $\mathcal{G}_{j}$, proven to be the linear combination that maximizes the sum of squared correlations with the alternatives located in $\mathcal{G}_{j}$ (Vigneau and Qannari, 2003). The loading $k_{ij}$ measures the correlation between the alternative $i$ and $PC_{\mathcal{G}_{j}}$. (ii) Select the alternative with the largest loading $k_{ij}$ as the representative "prototype" of $\mathcal{G}_{j}$ (Al-Kandari and Jolliffe, 2001): $\tau_{j}=\mathop{\arg\max}_{i\in\mathcal{G}_{j}}k_{ij}.$
Moreover, as shown in Proposition 1, bounding the number of alternatives within each cluster by $p_{m}$ is essential for asymptotic sample optimality. Fortunately, the hierarchical structure of the $\mathcal{A}\mathcal{C}$ and $\mathcal{A}\mathcal{C}^{+}$ algorithms makes this easy to enforce. We set an upper limit $p_{m}$ on cluster size—once reached, additional alternatives are assigned to the second closest cluster. This also provides a practical guideline for choosing $k$, which can be set to approximately 2–3 times $\frac{p}{p_{m}}$. In practice, the choice of $k$ is flexible, as the exact number of clusters is not critical so long as highly correlated alternatives are grouped together. In addition, in the supplementary material, we describe two additional steps to address potential issues: (i) unbalanced cluster sizes and (ii) the presence of alternatives with similar means within the same cluster.
Algorithm 2 Few-shot Alternative Clustering $\mathcal{A}\mathcal{C}^{+}$
Input: sample size $n$, the number of clusters $k$, $p_{s}$, $p_{q}$.
Step 1: Randomly split the set $\mathcal{P}$ into two sets, $\mathcal{P}_{s}$ and $\mathcal{P}_{q}$, with sizes $p_{s}$ and $p_{q}$, respectively.
Step 2 (Selecting Prototypes): On a single processor, simulate $n$ samples for each alternative in $\mathcal{P}_{s}$ and calculate the estimated covariance matrix $\hat{\mathrm{\Sigma}}_{n}^{\mathcal{P}_{s}}$. Input $\hat{\mathrm{\Sigma}}_{n}^{\mathcal{P}_{s}}$ into the algorithm $\mathcal{A}\mathcal{C}$ and output the clustering result $\Pi_{n}^{\mathcal{P}_{s}}$. Within $\mathcal{P}_{s}$, calculate the prototype $\tau_{j}=\mathop{\arg\max}_{i\in\mathcal{G}_{j}}k_{ij}$ of cluster $\mathcal{G}_{j}$, $\forall j\in\{1,2,...,k\}$. Let $\mathcal{P}_{p}=\{\tau_{1},\cdots,\tau_{k}\}$.
Step 3 (Matching): Send a copy of $\mathcal{P}_{p}$ to each processor. Randomly and equally allocate $\mathcal{P}_{q}$ to each processor. For any $j\in\mathcal{P}_{q}$, simulate alternative $j$ for $n$ times and calculate $\hat{r}_{\tau_{i},j}^{n}$ for any $i\in\{1,2,...,k\}$. $G_{n}(j)=\arg\max\limits_{i\in\{1,2,...,k\}}{\hat{r}_{\tau_{i},j}^{n}}$.
 Return the partition $\Pi_{n}$.
5.2 Computation of $\mathrm{PCC}$
Next, we establish a computable lower bound for $\mathrm{PCC}_{\mathcal{A}\mathcal{C}^{+}}$, the statistical guarantee of $\mathcal{A}\mathcal{C}^{+}$. This computable bound is practically useful for guiding the initialization sample size required for clustering. Let $\Gamma\triangleq\{(ab,ac)|G(a)=G(b),G(a)\neq G(c),a,b,c\in\mathcal{P},a\neq b\}$ be the collection of pairs of overlapping intra-cluster and inter-cluster correlations. $\Gamma_{s}\triangleq\{(ab,ac)\in\Gamma|a,b,c\in\mathcal{P}_{s}\}$ and $\Gamma_{q}\triangleq\{(a\tau_{i},a\tau_{j})\in\Gamma|a\in\mathcal{P}_{q},i=G(a),j\in\{1,2,\dots k\}\setminus\{i\}\}$ are two subsets of $\Gamma$. Fisher’s $z$ transformation of correlation coefficients, defined as $z\left(r\right)\triangleq\frac{1}{2}ln\frac{1+r}{1-r}$, is monotonically increasing in $(0,1)$. Therefore, Assumption 3 implies that $z(r_{ab})>z(r_{ac})$ for any pair $(ab,ac)\in\Gamma$. Borrowing the idea of indifference zone in R&S literature, in the following proposition we introduce a correlation indifference parameter $\delta_{c}>0$ and assume $z(r_{ab})>z(r_{ac})+\delta_{c}$.
Proposition 2
If Assumption 3 holds and all clusters have equal sizes, and the $\mathcal{A}\mathcal{C}^{+}$ algorithm uses $N$ samples to perform clustering, then
$\mathrm{PCC}_{\mathcal{A}\mathcal{C}^{+}}\geq\big(1-k(1-1/k)^{p_{s}}\big)P\bigg(\bigcap_{(ab,ac)\in\Gamma_{s}}\{\hat{r}_{ab}^{n}>\hat{r}_{ac}^{n}\}\bigg)P\bigg(\bigcap_{(ab,ac)\in\Gamma_{q}}\{\hat{r}_{ab}^{n}>\hat{r}_{ac}^{n}\}\bigg).$ (5)
Let $\Gamma_{\star}$ be either $\Gamma_{s}$ or $\Gamma_{q}$. If $z(r_{ab})>z(r_{ac})+\delta_{c}$ for any $(ab,ac)\in\Gamma_{\star}$, then
$P\bigg(\bigcap_{(ab,ac)\in\Gamma_{\star}}\{\hat{r}_{ab}^{n}>\hat{r}_{ac}^{n}\}\bigg)\geq\sum_{(ab,ac)\in\Gamma_{\star}}{\Phi\Bigg({\delta_{c}}{\sqrt{\frac{n-3}{2(1-\hat{r}_{bc}^{n})h(a,b,c)}}}\Bigg)}-(\left\lvert\Gamma_{\star}\right\rvert-1),$ (6)
where $h(a,b,c)=\frac{1-f\cdot\bar{R}^{2}}{1-{\bar{R}}^{2}}$, with $\bar{R}^{2}(a,b,c)=\frac{\left(\hat{r}_{ab}^{n}\right)^{2}+\left(\hat{r}_{bc}^{n}\right)^{2}}{2}$ and $f(a,b,c)=\frac{1-\hat{r}_{bc}^{n}}{2\left(1-\bar{R}^{2}\right)}$.
Additionally, the computation of $\mathrm{PCC}_{\mathcal{A}\mathcal{C}^{+}}$ when cluster sizes are unequal can be found in the supplementary material. As $n\to\infty$, the lower bound of $\mathrm{PCC}_{\mathcal{A}\mathcal{C}^{+}}$ converges to $1-k(1-\frac{1}{k})^{p_{s}}$. This bound approaches 1 as $p_{s}$ increases. P3C enjoys the advantage of not requiring an excessively high PCC or accurate correlation estimation. As long as the accuracy surpasses that of random clustering, an improvement in sample efficiency is ensured. Moreover, with Proposition 2, one can determine the required sample size to achieve a given clustering precision (see the supplementary material).
6 Numerical Experiments
In Section 6.1, we present a simple example to illustrate the key theoretical result of this paper: the mean-correlation interaction (Theorem 2, Corollary 1, and Lemma 1). Section 6.2 evaluates P3C under fixed-precision constraints, while Section 6.3 examines its performance under fixed-budget constraints.
6.1 Illustrative Example
Consider a group of 5 alternatives: $(X_{1},X_{2},X_{3},X_{4},X_{5})$, each following a normal distribution. The mean and covariance parameters are given by $\bm{\mu}=(\mu_{1},\mu_{2},\mu_{3},\mu_{4},\mu_{5})=(2.1,2.0,1.95,1.9,1.9)$ and
$\Sigma=\left(\begin{array}[]{ccccc}0.1&x&x&x&x\\ x&0.1&0.01&0.01&y\\ x&0.01&0.1&0.01&y\\ x&0.01&0.01&0.1&y\\ x&y&y&y&0.1\end{array}\right).$
All alternatives share the same variance. Alternative 1 has the highest mean, and its covariance with other alternatives is $x$ ($0\leq x<0.1$). Alternative 5 is among the worst performers, and its covariance with alternatives 2, 3 and 4 is $y$ ($0\leq y<0.1$). To illustrate the proposed mean-correlation interaction theory (Theorem 2, Corollary 1 and Lemma 1), we calculate $\text{PCS}(1)$ and $\text{PCS}(5)$ with different $x$ and $y$, keeping the mean and variance parameters constant. The sample size for each alternative is 10 (total sample size $N=50$). Table 1 presents the values of $\text{PCS}(1)$ and $\text{PCS}(5)$ when $x$ takes on the values of 0.01, 0.02, 0.03, and 0.05, and $y$ takes on the values of 0, 0.02, 0.04, and 0.06, respectively (there is no value in the table for "x=0.05,y=0.06x=0.05,y=0.06" because $\Sigma$ is not positive definite). The actual values of individual PCS are calculated using Monte Carlo numerical integration. The experimental results align with our theory of mean-correlation interaction: increasing correlation promotes alternatives with high means while suppressing those with low means. When $y$ is held constant, $\text{PCS}(1)$ increases as $x$ increases, consistent with Theorem 2 and Corollary 1. When $x$ is fixed, $\text{PCS}(1)$ also increases with $y$, where $y$ represents correlations between alternative 5 and alternatives 2-4, not directly associated with alternative 1. This result is consistent with Lemma 1. Conversely, $\text{PCS}(5)$ decreases as $x$ (or $y$) increases, as predicted by Theorem 2.
Table 1: $\text{PCS}(1)$ and $\text{PCS}(5)$ with Different $x$ and $y$ ($N=50$).
$y=0$ $y=0.02$ $y=0.04$ $y=0.06$
$\text{PCS}(1)$ $\text{PCS}(5)$ $\text{PCS}(1)$ $\text{PCS}(5)$ $\text{PCS}(1)$ $\text{PCS}(5)$ $\text{PCS}(1)$ $\text{PCS}(5)$
$x=0.01$ 0.6707 0.0349 0.6741 0.0269 0.6875 0.1777 0.6879 0.0051
$x=0.02$ 0.6852 0.0331 0.6904 0.0255 0.7003 0.0168 0.7018 0.0050
$x=0.03$ 0.6979 0.0288 0.7141 0.0233 0.7171 0.0154 0.7274 0.0048
$x=0.05$ 0.7506 0.0197 0.7662 0.0166 0.7690 0.0119
6.2 Fixed-precision R&S: Drug Discovery
We test the performance of P3C on the narcotic analgesics drug discovery problem introduced in Negoescu et al. (2011), which studies a set of $11\times 8\times 5\times 6\times 11\times 3\approx 10^{5}$ Benzomorphans, each representing a potential alternative drug. Simulation data are generated based on the Free-Wilson model (Free and Wilson, 1964): each drug’s value is modeled as the sum of the values of its constituent atomic groups. The value of each atomic group is treated as an independent normally distributed random variable, with mean estimated via regression on experimental data from Katz et al. (1977), and variance randomly drawn from $N(0,0.1)$, excluding negative values. Since the Free-Wilson model inherently uses CRN, structurally similar drugs tend to be highly correlated. We adopt the fixed-precision formulation of R&S. Our objective is to identify the best drug while ensuring $\mathrm{PCS}_{\mathrm{trad}}>0.9$. In this experiment, we compare P3C-KN and P3C-KT with the traditional stage-wise procedure from Rinott (1978), the Good Selection Procedure (GSP) in Ni et al. (2017), and the standard KT and KN without P3C. The indifference-zone parameter is set to $\delta=0.1$, with $p_{m}=1000$ and $k=104$. The initialization sample size is set to $N_{0}=20$, and both correlation estimation and clustering in P3C are performed solely using these initialization samples, without requiring any additional sampling. Experiments are conducted on a commercial cloud platform using a computing cluster with 104 processors. We adopt MATLAB’s Parallel Computing Toolbox to manage workload distribution and inter-processor communication. For fairness, Rinott, GSP, standard KT, and standard KN are all implemented in the same parallel environment as P3C, using the traditional divide and conquer strategy and CRN. Each experiment is repeated 20 times, and results are averaged.
The left panel of Figure 3 reports the total sample sizes required by P3C-KT, P3C-KN, KT, KN, Rinott, and GSP across different values of $p$. P3C-KN and P3C-KT consistently outperform the other algorithms. Both algorithms exhibit linear growth in sample size with respect to $p$, confirming the $\mathcal{O}(p)$ sample complexity predicted by Proposition 1. On average, P3C-KN requires only 46% of the sample size used by standard KN, while P3C-KT uses 68% of that of KT. These results support the analysis in Section 4, which shows how correlation-based clustering reduces the sample complexity in parallel R&S. Notably, the slope of KT decreases visibly after applying P3C, reflecting a reduction in the $\mathcal{O}(p)$ coefficient, consistent with Theorem 3. In contrast, the classical stage-wise Rinott procedure demands 30–50 times more samples than the other fully sequential methods. The right panel of Figure 3 reports the wall-clock time of the experiments. Compared to the standard "divide and conquer" framework, P3C introduces only one additional clustering step, making clustering the primary source of extra overhead. The table includes clustering time using the $\mathcal{AC}^{+}$ algorithm, where $p_{s}$ is set to one-tenth of the total alternatives. Even when $p$ reaches $2^{16}$, clustering takes only 11.714 seconds, which is negligible relative to the total R&S runtime. Notably, KT-P3C and KN-P3C are significantly faster than their non-P3C counterparts. This is because the reduction in sample complexity achieved by P3C far outweighs the clustering cost.
$p$ $2^{11}$ $2^{13}$ $2^{15}$ $2^{16}$
Clustering Time 0.086 0.208 2.438 11.714
KT 2.213 9.379 47.708 129.790
P3C-KT 1.288 3.306 18.188 48.830
KN 3.708 12.057 65.770 168.114
P3C-KN 8.123 17.803 37.573 89.200
GSP 2.646 6.686 71.556 278.306
Figure 3: Comparison of computational time and required sample size under different numbers of alternatives pp. The left panel shows the required total sample size ($\times 10^{4}$) as a function of $p$, while the right panel presents the corresponding wall clock time (seconds).
According to Theorem 3, $\mathrm{PCC}$ significantly influences the performance of P3C. Using the $\mathcal{A}\mathcal{C}^{+}$ algorithm, we investigate $\mathrm{PCC}_{\mathcal{A}\mathcal{C}^{+}}$ under different sample sizes and different sizes of the support set ($p_{s}=25,50,75,100,150$). We focus on the first 1024 drugs and set the cluster number to $k=8$. The results are shown in Figure 4. As the sample size approaches infinity, consistent with Proposition 2, $\mathrm{PCC}_{\mathcal{A}\mathcal{C}^{+}}$ converges to $1-k(1-1/k)^{p_{s}}$ (denoted as $P(D)$), which is strictly less than 1. Larger $p_{s}$ values lead to a larger $P(D)$. When $p_{s}=25$, the PCC is notably constrained by the $P(D)$ term and cannot exceed 0.3. As $p_{s}$ increases to 50, $\mathrm{PCC}$ approaches acceptably close to 1.
Figure 4: Comparison of $\mathrm{PCC}_{\mathcal{A}\mathcal{C}^{+}}$ with different support set sizes $p_{s}$.
6.3 Fixed-budget R&S: Neural Architecture Search
Neural Architecture Search (NAS), a key challenge in deep learning, aims to identify the best-performing neural network architecture for a specific task, given that evaluating a neural architecture on a test dataset is computationally expensive. The general NAS problem can be divided into two phases: architecture search (including the search space and search strategy) and architecture performance evaluation. Suppose that we have already obtained a set $\mathcal{P}$ of $p$ alternative architectures through certain search methods. The best architecture is defined as the one that maximizes generalization accuracy: $1]\triangleq\mathop{\arg\max}_{i\in\mathcal{P}}\mathbb{E}(\text{ACC}_{i})$, where $\text{ACC}_{i}$ denotes the accuracy of alternative $i$. The expectation is theoretically taken over the probability space of all unseen data points, which is impossible to compute exactly, so we estimate it using Monte Carlo simulation over the test dataset. Therefore, NAS can be viewed as a discrete simulation optimization problem, where the simulation model is a neural network. Testing one architecture on a batch of test data corresponds to one simulation run, and the resulting accuracy is treated as the simulation output. Given the limited size of the test dataset, which corresponds to a finite amount of simulation resources, we formulate NAS as a fixed-budget R&S problem. We compare the following algorithms: OCBA; CBA (correlated budget allocation) from Fu et al. (2007), which extends OCBA by accounting for correlations between alternatives; OCBA and CBA combined with P3C, denoted as P3C-OCBA and P3C-CBA; FBKT from Hong et al. (2022); and the naive equal allocation (EA). Additionally, standard OCBA, CBA, FBKT, and EA are also implemented in a parallel computing environment using the traditional "divide and conquer” strategy, as described in Section 3.
NAS setting. We conduct experiments on the CIFAR-10 dataset, which consists of 50,000 training images and 10,000 test images for image classification. The architecture search phase is implemented in PyTorch using the state-of-the-art Single-Path One-Shot (SPOS) method (Guo et al., 2020). The search space is a single-path supernet composed of 20 choice blocks connected in series, each with 4 choices. We set the number of alternative architectures to $p=10^{5}$.
R&S setting. The R&S phase is conducted under the same parallel computing environment as described in Section 6.2. After obtaining $10^{5}$ alternative architectures, we use P3C-OCBA, OCBA, P3C-CBA, CBA, FBKT and EA to allocate computational resources for performance evaluation and select the best architecture. The R&S performance is evaluated using $\text{PCS}_{\text{trad}}$. One simulation observation corresponds to testing an architecture on a batch of 32 images. Therefore, the maximum sample size for each alternative is $\frac{10000}{32}\approx 312$. The initialization sample size is set to $N_{0}=20$. The $\text{PCS}_{\text{trad}}$ values are directly estimated based on 1000 independent macro replications. The true best $[1]$ (invisible to the users and used solely for the final $\text{PCS}_{\text{trad}}$ calculation) is estimated by selecting the alternative with the highest accuracy on the full test set. Moreover, in the Stage 2 of "divide and conquer" and P3C, the $\text{PCS}_{\text{trad}}$ values vary across different processors. The overall $\text{PCS}_{\text{trad}}$ for Stage 2 is approximated by a weighted average based on the number of alternatives on each processor.
Total Sample Size 1.0 1.5 2.0 2.5 3.0
Clustering Time 10.2 10.0 10.1 10.0 10.2
CBA 29.9 31.2 33.7 35.6 36.7
P3C-CBA 42.1 43.4 46.4 46.2 47.1
OCBA 29.0 29.4 33.0 34.7 35.8
P3C-OCBA 39.1 39.2 43.0 45.2 46.0
FBKT 134.1 135.2 135.4 138.2 140.9
Figure 5: Comparison of different R&S procedures for different total sample sizes. The left panel shows the $\text{PCS}_{\text{trad}}$ metric comparison, while the right panel presents the corresponding wall clock time (seconds) for different total sample sizes ($\times 10^{7}$).
Figure 5 presents the experimental results under different total sample sizes. This is a large-scale R&S problem with limited simulation resources. As shown in the left panel, $\text{PCS}_{\text{trad}}$ remains extremely low under classic strategies such as OCBA and CBA. FBKT performs slightly better but still remains below 0.05. However, with P3C, both CBA and OCBA improve by more than an order of magnitude and surpass FBKT, demonstrating the effectiveness of P3C in large-scale fixed-budget R&S. Notably, P3C-CBA performs slightly better than P3C-OCBA by leveraging correlation information in the sampling policy. The right panel shows the wall-clock time for different algorithms (excluding the lengthy evaluation on the test data and reporting only the time spent on R&S). Since the total number of alternatives is fixed, the clustering time remains around 10 seconds. The additional time for P3C-CBA and P3C-OCBA, compared to their non-P3C counterparts, is primarily due to the clustering step. FBKT, a dynamic procedure requiring multiple rounds of comparisons, incurs significantly higher computational time.
7 Concluding Remarks
In this paper, both theoretical and empirical results indicate that leveraging correlation is an effective way to improve sample efficiency in large-scale R&S. The P3C procedure presented in this paper is a ready-to-use tool that achieves significant improvements without requiring excessively high clustering accuracy; it only needs to identify which alternatives are highly correlated.
[…]
Appendices
Appendix A The Gradient Analysis with respect to Correlation Information
A.1 Proofs of Theorem 2, Corollary 1 and Lemma 1.
Here, we prove a more general result that encompasses not only the derivative of PCS with respect to the correlation information but also the derivative with respect to variance information.
(a) $\frac{\partial PCS\left(\tau\right)}{\partial\sigma_{\tau}}=D^{+}+D^{-}+I$, where $D^{+}=\sum_{i\in\mathcal{I}^{+}}D^{\tau}_{i}$, $D^{-}=\sum_{i\in\mathcal{I}^{-}}D_{i}^{\tau}$, and the term $D^{\tau}_{i}>0$ for $i\in\mathcal{I}^{+}(\tau)$, $D^{\tau}_{i}<0$ for $i\in\mathcal{I}^{-}(\tau)$;
(b) Theorem 2: $\forall i\in\mathcal{P}\setminus\{\tau\}$, $\frac{\partial PCS\left(\tau\right)}{\partial r_{\tau i}}={\widetilde{D}}^{\tau}_{i}+{\widetilde{I}}^{\tau}_{i}$, where ${\widetilde{D}}^{\tau}_{i}>0$ for $i\in\mathcal{I}^{-}(\tau)$ and <0<0 for $i\in\mathcal{I}^{+}(\tau)$.
${\widetilde{I}}^{\tau}_{i}=\sum_{j\in\{1,\cdots,p\}\setminus\{i,\tau\}}{\widetilde{I}}^{\tau,j}_{i}$. $\mathrm{sign}({\widetilde{I}}^{\tau,j}_{i})=\mathrm{sign}\left(-\sigma_{i}^{2}(N_{i})^{-1}+{r_{\tau,i}\sigma_{\tau}\sigma_{i}}{(N_{\tau,i})^{-1}}-{r_{\tau,j}\sigma_{\tau}\sigma_{j}}{(N_{\tau,j})^{-1}}+{r_{i,j}\sigma_{i}\sigma_{j}}{(N_{i,j})^{-1}}\right)$;
(c) Lemma 1: $\forall i,j\in\mathcal{P}\setminus\{\tau\}$, $\frac{\partial\text{PCS}\left(\tau\right)}{\partial r_{ij}}\geq 0$. If $\mathscr{D}^{\tau}_{i}<\mathscr{I}^{\tau}_{i}$, then $\frac{\partial\text{PCS}\left(\tau\right)}{\partial r_{ij}}=o\bigg(|\frac{\partial\text{PCS}\left(\tau\right)}{\partial r_{\tau i}}|\bigg)$.
The notations in the appendix differ slightly from those in the main text. The symbols $D^{\tau}_{i}$ and $I^{\tau}_{i}$ in the main text correspond to $\widetilde{D}^{\tau}_{i}$ and $\widetilde{I}^{\tau}_{i}$ here. The proof of Corollary 1 is embedded in (b). Moreover, it is important to note that, in this subsection, when it comes to asymptotic results, we consider the asymptotic regime where sample size $N$ goes to infinity while keeping $p$ fixed. If Assumption 2 holds and $p$ is fixed, $\lambda^{\tau}_{i},\lambda^{\tau}_{j}=\mathcal{O}(\frac{1}{N})$, and then we have $\tilde{r}^{\tau}_{i,j}=\mathcal{O}(1),$ meaning that its order does not increase with the growth of $N$. Let $f_{y^{\tau}}$ denote the density of $\bm{y^{\tau}}=(y^{\tau}_{1},...,y^{\tau}_{\tau-1},y^{\tau}_{\tau+1},...,y^{\tau}_{p})\sim N(0,\Phi^{\tau})$. {pf} (a) The derivative of $\text{PCS}\left(\tau\right)$ with respect to the variance of $\tau$ is given by
$\frac{\partial\text{PCS}\left(\tau\right)}{\partial\sigma_{\tau}}=\frac{\partial P\left(y_{1}{}^{\tau}>-d_{1}{}^{\tau},\ldots,y_{p}{}^{\tau}>-d_{p}{}^{\tau}\right)}{\partial\sigma_{\tau}}=\sum_{i=1,\cdots,p,\ i\neq\tau}\frac{\partial d_{i}{}^{\tau}}{\partial\sigma_{\tau}}\frac{\partial\mathrm{PCS(\tau)}}{\partial d_{i}^{\tau}}+2\sum_{1\leq i<j\leq p,\ i,j\neq\tau}\frac{\partial\text{PCS}(\tau)}{\partial{{\widetilde{r}}_{i,j}}^{\tau}}\frac{\partial\widetilde{r}_{i,j}^{\tau}}{\partial\sigma_{\tau}}$ (7)
$=\frac{\partial d_{1}{}^{\tau}}{\partial\sigma_{\tau}}\int_{-d^{\tau}_{2}}^{\infty}\cdots\int_{-d_{p}{}^{\tau}}^{\infty}f_{y^{\tau}}\left(-d_{1}{}^{\tau},y_{2},\ldots,y_{p}\right)dy_{2}\cdots dy_{p}+\cdots+$
$\frac{\partial d_{p}{}^{\tau}}{\partial\sigma_{\tau}}\int_{-d_{1}{}^{\tau}}^{\infty}\cdots\int^{\infty}_{-d_{p-1}{}^{\tau}}f_{y^{\tau}}\left(y_{1},y_{2},\ldots,-d_{p}{}^{\tau}\right)dy_{1}\cdots dy_{p-1}+2\sum_{1\leq i<j\leq p,\ i,j\neq\tau}\frac{\partial\text{PCS}(\tau)}{\partial{{\widetilde{r}}_{i,j}}^{\tau}}\frac{\partial\widetilde{r}_{i,j}^{\tau}}{\partial\sigma_{\tau}}=D^{+}-D^{-}+I,$
where $D^{+}\triangleq\sum_{i\in\mathcal{P}\setminus\{\tau\}}\operatorname{max}\{0,D_{i}^{\tau}\}$$D^{-}\triangleq\sum_{i\in\mathcal{P}\setminus\{\tau\}}\operatorname{min}\{0,D_{i}^{\tau}\}$, $I=2\sum_{1\leq i<j\leq p,\ i,j\neq\tau}\frac{\partial\text{PCS}(\tau)}{\partial{{\widetilde{r}}_{i,j}}^{\tau}}\frac{\partial\widetilde{r}_{i,j}^{\tau}}{\partial\sigma_{\tau}}$, and
$D_{i}^{\tau}$ $\triangleq\frac{\partial d_{i}{}^{\tau}}{\partial\sigma_{\tau}}\int_{-d_{1}^{\tau}}^{\infty}\cdots\int_{-d_{i-1}^{\tau}}^{\infty}\int_{-d^{\tau}_{i+1}}^{\infty}\cdots\int_{-d_{\tau-1}^{\tau}}^{\infty}\int_{-d^{\tau}_{\tau+1}}^{\infty}\cdots\int_{-d_{p}{}^{\tau}}^{\infty}f_{y^{\tau}}\left(y_{1},\cdots,-d_{i}{}^{\tau},\cdots,y_{p}\right)dy_{1}\cdots dy_{p},$ (8)
$D^{+}$ is the sum of the positive parts of $D^{\tau}_{i}$ ($i\in\mathcal{P}\setminus\{\tau\}$) and $D^{-}$ is the sum of negative parts of them. Notice that $d_{i}^{\tau}=\frac{\mu_{\tau}-\mu_{i}}{\sqrt{\lambda_{i}^{\tau}}}$ is not necessarily non-negative as $\mu_{\tau}$ does not always exceed $\mu_{i}$. If $i\in\mathcal{I}^{+}(\tau)\triangleq\{i|\mu_{i}>\mu_{\tau},i\in\mathcal{P}\setminus\{\tau\}\}$, $\frac{\partial d_{i}{}^{\tau}}{\partial\sigma_{\tau}}>0$, the term $D_{i}>0$, which will only contribute to $D^{+}$. Conversely, if $i\in\mathcal{I}^{-}$, the term $D_{i}<0$ and it is only accounted for in $D^{-}$. This means that $D^{+}=\sum_{i\in\mathcal{I}^{+}}D_{i}$ and $D^{-}=\sum_{i\in\mathcal{I}^{-}}\left|D_{i}\right|$.
In the following part, we will evaluate the growth rate of the terms $D_{i}$ and $I$ as the total sample size $N$ increases. First, we prove that $\frac{\partial\text{PCS}(\tau)}{\partial{{\widetilde{r}}_{i,j}}^{\tau}}$ is bounded. With Slepian normal comparison lemma Azaïs and Wschebor (2009), we have 0≤PCS(τ;~ri,jτ+δ)-PCS(τ;~ri,jτ)δ≤arcsin(~ri,jτ+δ)-arcsin(~ri,jτ)2πδexp(-(diτ)2+(djτ)22(1+max{|~ri,jτ+δ|,|~ri,jτ|})).
Let $\delta\rightarrow 0$, then
$0\leq\frac{\partial\text{PCS}(\tau)}{\partial{{\widetilde{r}}_{i,j}}^{\tau}}\leq\frac{1}{2\pi\sqrt{1-({{\widetilde{r}}_{i,j}}^{\tau})^{2}}}\exp{\bigg(-\frac{(d_{i}^{\tau})^{2}+(d_{j}^{\tau})^{2}}{2(1+|{{\widetilde{r}}_{i,j}}^{\tau}|)}\bigg)}.$ (9)
According to Assumption 2, we have $N_{i}=\mathcal{O}(N)$, $\forall i\in\mathcal{P}$ ($p$ is fixed). Then $d^{\tau}_{i}=\mathcal{O}(\sqrt{N})$ and ${\widetilde{r}}_{i,j}^{\tau}=\mathcal{O}(1)$. Then $\frac{\partial\text{PCS}(\tau)}{\partial{{\widetilde{r}}_{i,j}}^{\tau}}=\mathcal{O}\bigg(\exp{(-\frac{(d_{i}^{\tau})^{2}+(d_{j}^{\tau})^{2}}{2(1+|{{\widetilde{r}}_{i,j}}^{\tau}|)})}\bigg),$ and we can introduce a positive constant $\mathscr{I}^{\tau}_{ij}$ such that $\frac{\partial\text{PCS}(\tau)}{\partial{{\widetilde{r}}_{i,j}}^{\tau}}=\mathcal{O}(e^{-\mathscr{I}^{\tau}_{ij}N})$. We define $\mathscr{I}^{\tau}_{i}=\min_{j\neq i}{\mathscr{I}^{\tau}_{ij}},$ and $\mathscr{I}^{\tau}=\min_{i\in\mathcal{P}\setminus\{\tau\}}{\mathscr{I}^{\tau}_{i}}.$ It is straightforward to verify that $\frac{\partial{{\widetilde{r}}_{i,j}}^{\tau}}{\partial\sigma_{\tau}}=\mathcal{O}(1)$ by definition. Then we have $I=\mathcal{O}\bigg(\exp{\big(-\min_{i,j\in\mathcal{P}\setminus\{\tau\}}\frac{(d_{i}^{\tau})^{2}+(d_{j}^{\tau})^{2}}{2(1+|{{\widetilde{r}}_{i,j}}^{\tau}|)}\big)}\bigg)=\mathcal{O}(e^{-\mathscr{I}^{\tau}N}).$ By definition, we have $\frac{\partial d^{\tau}_{i}}{\partial\sigma_{\tau}}=\mathcal{O}(\sqrt{N}).$ Next, we will evaluate the following integral for each $i\in\mathcal{P}\setminus\{\tau\}$:
$\frac{D_{i}^{\tau}}{\frac{\partial d_{i}{}^{\tau}}{\partial\sigma_{\tau}}}$ $=\int_{-d_{1}^{\tau}}^{\infty}\cdots\int_{-d_{i-1}^{\tau}}^{\infty}\int_{-d^{\tau}_{i+1}}^{\infty}\cdots\int_{-d_{\tau-1}^{\tau}}^{\infty}\int_{-d^{\tau}_{\tau+1}}^{\infty}\cdots\int_{-d_{p}{}^{\tau}}^{\infty}f_{y^{\tau}}\left(y_{1},\cdots,-d_{i}{}^{\tau},\cdots,y_{p}\right)dy_{1}\cdots dy_{p}$ (10)
$=f_{y^{\tau}_{i}}(-d_{i}^{\tau})\int_{-d_{1}^{\tau}}^{\infty}\cdots\int_{-d_{i-1}^{\tau}}^{\infty}\int_{-d^{\tau}_{i+1}}^{\infty}\cdots\int_{-d_{\tau-1}^{\tau}}^{\infty}\int_{-d^{\tau}_{\tau+1}}^{\infty}\cdots\int_{-d_{p}{}^{\tau}}^{\infty}f_{y^{\tau}|y^{\tau}_{i}}\left(y_{1},\cdots,y_{p}\right)dy_{1}\cdots dy_{p}.$
$f_{y^{\tau}_{i}}$ is the marginal density of $y^{\tau}_{i}$, which is standard normal, and $f_{y^{\tau}|y^{\tau}_{i}}$ is the conditional density of $y^{\tau}$ given $y^{\tau}_{i}$, which is a multivariate normal distribution $N(\widetilde{\mu}_{i},\widetilde{\Sigma}_{i})$ of dimension p2p-2 with mean and covariance given by $\widetilde{\mu}_{i}=(-d_{i}^{\tau}\widetilde{r}^{\tau}_{i,1},\cdots,-d_{i}^{\tau}\widetilde{r}^{\tau}_{i,p})$ and
$\widetilde{\Sigma}_{i}=\left(\begin{array}[]{cccc}1-(\widetilde{r}^{\tau}_{i,1})^{2}&\ \widetilde{r}^{\tau}_{1,2}-\widetilde{r}^{\tau}_{i,1}\widetilde{r}^{\tau}_{i,2}&\ \cdots&\ \widetilde{r}^{\tau}_{1,p}-\widetilde{r}^{\tau}_{i,1}\widetilde{r}^{\tau}_{i,p}\\ \widetilde{r}^{\tau}_{2,1}-\widetilde{r}^{\tau}_{i,2}\widetilde{r}^{\tau}_{i,1}&\ 1-(\widetilde{r}^{\tau}_{i,2})^{2}&\ \cdots&\ \widetilde{r}^{\tau}_{2,p}-\widetilde{r}^{\tau}_{i,2}\widetilde{r}^{\tau}_{i,p}\\ \vdots&\vdots&\ddots&\vdots\\ \widetilde{r}^{\tau}_{p,1}-\widetilde{r}^{\tau}_{i,p}\widetilde{r}^{\tau}_{i,1}&\cdots&\cdots&1-(\widetilde{r}^{\tau}_{i,p})^{2}\end{array}\right)_{(p-2)\times(p-2)}.$ (11)
By transforming variables, we have
$\int_{-d_{1}^{\tau}}^{\infty}\cdots\int_{-d_{p}^{\tau}}^{\infty}f_{y^{\tau}|y^{\tau}_{i}=-d_{i}^{\tau}}\left(y_{1},\cdots,y_{p}\right)dy_{1}\cdots dy_{p}=\int_{-\widetilde{d}_{1}^{\tau}}^{\infty}\cdots\int_{-\widetilde{d}_{p}^{\tau}}^{\infty}f_{Z}\left(z_{1},\cdots,z_{p}\right)dz_{1}\cdots dz_{p},$ (12)
where $\widetilde{d}_{j}^{\tau}=-d_{j}^{\tau}+d_{i}^{\tau}\widetilde{r}^{\tau}_{i,j}$ and $f_{Z}$ is the density of a multivariate normal distribution with zero mean and covariance matrix $\Sigma^{Z}=\widetilde{\Sigma}_{i}$. Then, the integral (LABEL:transform) is rewritten as
$P_{Z}(z_{j}>-\widetilde{d}_{j}^{\tau},j\in\mathcal{P}\setminus\{i,\tau\})=P_{Z}(\{z_{r}>-\widetilde{d}_{r}^{\tau},r\in\mathcal{S}_{-i}^{+}(\tau)\}\cap\{z_{s}>-\widetilde{d}_{s}^{\tau},s\in\mathcal{S}_{-i}^{-}(\tau)\})$ (13)
$\leq\min\{P_{Z}(z_{r}>-\widetilde{d}_{r}^{\tau},r\in\mathcal{S}_{-i}^{+}(\tau)),P_{Z}(z_{s}>-\widetilde{d}_{s}^{\tau},s\in\mathcal{S}_{-i}^{-}(\tau))\},$
where $\mathcal{S}_{-i}^{+}(\tau)\triangleq\{r\in\mathcal{P}\setminus\{i,\tau\}|-\widetilde{d}_{r}^{\tau}>0\}$ and $\mathcal{S}_{-i}^{-}(\tau)\triangleq\{s\in\mathcal{P}\setminus\{i,\tau\}|-\widetilde{d}_{s}^{\tau}<0\}$.
Next we evaluate the two terms within the "min" operator respectively. The term $P_{Z}(z_{s}>-\widetilde{d}_{s}^{\tau},s\in\mathcal{S}_{-i}^{-}(\tau))\geq P_{Z}(z_{s}>0,s\in\mathcal{S}_{-i}^{-}(\tau))$ is lower bounded by a positive value that will not reduce to 0 as $N$ increases. However, the other term $P_{Z}(z_{r}>-\widetilde{d}_{r}^{\tau},r\in\mathcal{S}_{-i}^{+}(\tau))$ will reduce to 0 as $N$ grows since $d_{r}^{\tau}=\mathcal{O}(\sqrt{N})$. Next we evaluate the rate of convergence to 0. With Hashorva and Hüsler (2003), there exists a subset $S$ of $\mathcal{S}_{-i}^{+}(\tau)$ such that
$P_{Z}(z_{r}>-\widetilde{d}_{r}^{\tau},r\in\mathcal{S}_{-i}^{+}(\tau))=\mathcal{O}\bigg(\mathrm{exp}(-\frac{Q^{\tau}_{S}}{2})\prod_{r\in S}h_{r}^{-1}\bigg),$ (14)
where $Q^{\tau}_{S}\triangleq\langle\bm{d}_{S},(\Sigma^{Z}_{S})^{-1}\bm{d}_{S}\rangle=\min_{\bm{x}\geq\bm{d}_{\mathcal{S}_{-i}^{+}(\tau)}}\langle\bm{x},(\Sigma^{Z}_{\mathcal{S}_{-i}^{+}(\tau)})^{-1}\bm{x}\rangle$, $\bm{x}=(x_{1},\cdots,x_{|\mathcal{S}_{-i}^{+}(\tau)|})\in\mathbb{R}^{|\mathcal{S}_{-i}^{+}(\tau)|}$, $\bm{d}=(-\widetilde{d}_{1}^{\tau},\cdots,-\widetilde{d}_{p}^{\tau})\in\mathbb{R}^{p-2}$ and $h_{r}=\mathcal{O}(\sqrt{N})$ is the $r$-th element of $(\Sigma^{Z}_{S})^{-1}\bm{d}_{S}$. For simplicity, we omit the polynomial term $\prod_{r\in S}h_{r}^{-1}$, as it does not affect the comparisons of exponentially decaying terms. Finally, with (10), (13) and (14), we have $D_{i}^{\tau}=\mathcal{O}\bigg(\frac{\partial d_{i}{}^{\tau}}{\partial\sigma_{\tau}}f_{y^{\tau}_{i}}(-d_{i}^{\tau})\mathrm{exp}(-\frac{\langle\bm{d}_{S},(\Sigma^{Z}_{S})^{-1}\bm{d}_{S}\rangle}{2})\bigg)=\mathcal{O}\bigg(\mathrm{exp}(-\frac{Q^{\tau}_{S}+(d_{i}^{\tau})^{2}}{2})\bigg).$ To simplify the form, we rewrite the quadratic form by introducing $\mathscr{D}^{\tau}_{i}>0$ such that $D_{i}^{\tau}=\mathcal{O}(e^{-\mathscr{D}^{\tau}_{i}N}).$
(b) Similar to (a), the derivative of $\text{PCS}\left(\tau\right)$ with respect to correlation information $\{r_{\tau,i}\}_{i\neq\tau}$ is given by
$\frac{\partial\text{PCS}\left(\tau\right)}{\partial r_{\tau i}}=\frac{\partial P\left(y_{1}{}^{\tau}>-d_{1}{}^{\tau},\ldots,y_{p}{}^{\tau}>-d_{p}{}^{\tau}\right)}{\partial r_{\tau i}}=\frac{\partial d_{i}{}^{\tau}}{\partial r_{\tau i}}\frac{\partial\mathrm{PCS(\tau)}}{\partial d_{i}^{\tau}}+2\sum_{1\leq i<j\leq p,\ i,j\neq\tau}\frac{\partial\text{PCS}(\tau)}{\partial{{\widetilde{r}}_{i,j}}^{\tau}}\frac{\partial\widetilde{r}_{i,j}^{\tau}}{\partial r_{\tau i}}$ (15)
$=\frac{\partial d_{i}{}^{\tau}}{\partial r_{\tau i}}\int_{-d_{1}^{\tau}}^{\infty}\cdots\int_{-d_{i-1}^{\tau}}^{\infty}\int_{-d_{i+1}^{\tau}}^{\infty}\cdots\int_{-d_{p}{}^{\tau}}^{\infty}f_{y^{\tau}}\left(y_{1},\cdots,-d_{i}{}^{\tau},\cdots,y_{p}\right)dy_{1}\cdots dy_{p}$
$+2\sum_{j\in\{1,\cdots,p\}\setminus\{i,\tau\}}\frac{\partial\text{PCS}(\tau)}{\partial{{\widetilde{r}}_{i,j}}^{\tau}}\frac{\partial\widetilde{r}_{i,j}^{\tau}}{\partial r_{\tau i}}={\widetilde{D}}^{\tau}_{i}+{\widetilde{I}}_{i},$
where
${\widetilde{D}}^{\tau}_{i}=\frac{\partial d_{i}{}^{\tau}}{\partial r_{\tau i}}\int_{-d_{1}^{\tau}}^{\infty}\cdots\int_{-d_{i-1}^{\tau}}^{\infty}\int_{-d_{i+1}^{\tau}}^{\infty}\cdots\int_{-d_{p}{}^{\tau}}^{\infty}f_{y^{\tau}}\left(y_{1},\cdots,-d_{i}{}^{\tau},\cdots,y_{p}\right)dy_{1}\cdots dy_{p},$ (16)
$\frac{\partial\widetilde{r}_{i,j}^{\tau}}{\partial r_{\tau i}}=\frac{\left(-\dfrac{\sigma^{2}_{i}}{N_{i}}+\dfrac{r_{\tau,i}\sigma_{\tau}\sigma_{i}}{N_{\tau i}}-\dfrac{r_{\tau,j}\sigma_{\tau}\sigma_{j}}{N_{\tau j}}+\dfrac{r_{i,j}\sigma_{i}\sigma_{j}}{N_{ij}}\right)\cdot\dfrac{\sigma_{\tau}\sigma_{i}}{N_{\tau i}}}{\left(\dfrac{\sigma^{2}_{\tau}}{N_{\tau}}+\dfrac{\sigma^{2}_{i}}{N_{i}}-2\dfrac{r_{\tau,i}\sigma_{\tau}\sigma_{i}}{N_{\tau i}}\right)^{3/2}\left(\dfrac{\sigma^{2}_{\tau}}{N_{\tau}}+\dfrac{\sigma^{2}_{j}}{N_{j}}-2\dfrac{r_{\tau,j}\sigma_{\tau}\sigma_{j}}{N_{\tau j}}\right)^{1/2}},$ (17)
${\widetilde{I}}_{i}=2\sum_{j\in\{1,\cdots,p\}\setminus\{i,\tau\}}\frac{\partial\text{PCS}(\tau)}{\partial{{\widetilde{r}}_{i,j}}^{\tau}}\frac{\partial\widetilde{r}_{i,j}^{\tau}}{\partial r_{\tau i}}$, $N_{\tau i}=\max\{N_{\tau},N_{i}\}$, $N_{\tau j}=\max\{N_{\tau},N_{j}\}$, and $N_{ij}=\max\{N_{i},N_{j}\}$. Similar to (a), we have ${\widetilde{D}}^{\tau}_{i}=\mathcal{O}\big(\exp\big(-\frac{Q^{\tau}_{S}+(d_{i}^{\tau})^{2}}{2}\big)\big)=\mathcal{O}(\sqrt{N}e^{-\mathscr{D}^{\tau}_{i}N}),$ and $\frac{\partial\widetilde{r}_{i,j}^{\tau}}{\partial r_{\tau i}}=\mathcal{O}(1),$ ${\widetilde{I}}_{i}=\mathcal{O}\bigg(\exp{\big(-\min_{j\neq i}\frac{(d_{i}^{\tau})^{2}+(d_{j}^{\tau})^{2}}{2(1+|{{\widetilde{r}}_{i,j}}^{\tau}|)}\big)}\bigg)=\mathcal{O}(e^{-\mathscr{I}^{\tau}_{i}N}),$ which concludes the proof of Corollary 1(a).
Next, we prove Corollary 1(b). If $\tau=[1]$, since $|\tilde{r}^{\tau}_{i,j}|<|\frac{d^{\tau}_{j}}{d^{\tau}_{i}}|$ according to the Assumption 1(a), the sign of $\widetilde{d}_{j}^{\tau}$ is the same as $d_{j}^{\tau}$. Then $\mathcal{S}_{-i}^{+}(\tau)=\{s\in\mathcal{P}\setminus\{i,\tau\}|-\widetilde{d}_{s}^{\tau}>0\}=\{s\in\mathcal{P}\setminus\{i,\tau\}|-d_{s}^{\tau}>0\}=\emptyset$, $\forall i\neq\tau$. Therefore, $Q^{\tau}_{S}=0$, and $\frac{Q^{\tau}_{S}+(d_{i}^{\tau})^{2}}{2}=\frac{(d_{i}^{\tau})^{2}}{2}<\min_{j\neq i}\frac{(d_{i}^{\tau})^{2}+(d_{j}^{\tau})^{2}}{2(1+|{{\widetilde{r}}_{i,j}}^{\tau}|)}$ with Assumption 1(b). Then $\mathscr{D}^{\tau}_{i}<\mathscr{I}^{\tau}_{i}$.
Next, we prove Corollary 1(c), which establishes the monotonic non-decreasing property of $\mathscr{D}^{\tau}_{i}(m)$ with respect to the ranking $m$ of $\tau$. We assume that only the ranking $m$ changes, while all other parameters (covariance matrix, sample sizes, and the absolute value of pairwise mean differences) remain fixed. To prove this, we only need to show that $\mathscr{D}^{\tau}_{i}(p)\geq\cdots\geq\mathscr{D}^{\tau}_{i}(2)\geq\mathscr{D}^{\tau}_{i}(1),\ \forall i\in\mathcal{P}\setminus\{\tau\}.$ As mentioned earlier, $\mathcal{S}_{-i}^{+}(\tau)=\{s\in\mathcal{P}\setminus\{i,\tau\}|-d_{s}^{\tau}>0\}$ due to Assumption 1(a), and it is straightforward to conclude that $\mathcal{S}_{-i}^{+}([1]])\subseteq\mathcal{S}_{-i}^{+}([2]])\subseteq\cdots\subseteq\mathcal{S}_{-i}^{+}([p]]),$ where $\mathcal{S}_{-i}^{+}([m]])$ denotes the corresponding $\mathcal{S}_{-i}^{+}(\tau)$ when $\tau=[m]$. This inclusion holds because, as the ranking $m$ of $\tau$ increases, more alternatives surpass $\tau$ in terms of their mean values. Then
$Q^{[p]}_{S}\geq\cdots\geq Q^{[2]}_{S}\geq Q^{[1]}_{S}$ (18)
can be proved by contradiction, where $Q^{[m]}_{S}$ denotes the corresponding $Q^{[\tau]}_{S}$ when $\tau=[m]$. Suppose that there exist $m<m^{\prime}$ such that $Q^{[m]}_{S}>Q^{[m^{\prime}]}_{S}$. Let x_∗∈R^|S_-i^+([m])|=argmin_xd_S_-i^+([m]) ⟨x,Σ^Z_S_-i^+([m])x ⟩ and x’_∗∈R^|S_-i^+([m’])|=argmin_xd_S_-i^+([m’]) ⟨x,Σ^Z_S_-i^+([m’])x ⟩. Then $\langle\bm{x}_{\ast},\Sigma^{Z}_{\mathcal{S}_{-i}^{+}([m])}\bm{x}_{\ast}\rangle>\langle\bm{x^{\prime}}_{\ast},\Sigma^{Z}_{\mathcal{S}_{-i}^{+}([m^{\prime}])}\bm{x^{\prime}}_{\ast}\rangle$. Since $\mathcal{S}_{-i}^{+}([m]])\subseteq\mathcal{S}_{-i}^{+}([m^{\prime}]])$, let $\tilde{\bm{x}}_{\ast}$ be the elements in $\bm{x^{\prime}}_{\ast}\in\mathbb{R}^{|\mathcal{S}_{-i}^{+}([m^{\prime}])|}$ with indices corresponding to $\mathcal{S}_{-i}^{+}([m]])$. Then we have $\langle\bm{x}_{\ast},\Sigma^{Z}_{\mathcal{S}_{-i}^{+}([m])}\bm{x}_{\ast}\rangle>\langle\bm{x^{\prime}}_{\ast},\Sigma^{Z}_{\mathcal{S}_{-i}^{+}([m^{\prime}])}\bm{x^{\prime}}_{\ast}\rangle\geq\langle\widetilde{\bm{x}}_{\ast},\Sigma^{Z}_{\mathcal{S}_{-i}^{+}([m])}\widetilde{\bm{x}}_{\ast}\rangle$, which contradicts to the optimality of $\bm{x}_{\ast}$. Therefore, (18) holds. Then according to the definition of $\mathscr{D}^{\tau}_{i}(m)$, we can conclude that $\mathscr{D}^{\tau}_{i}(p)\geq\cdots\geq\mathscr{D}^{\tau}_{i}(2)\geq\mathscr{D}^{\tau}_{i}(1),\ \forall i\in\mathcal{P}\setminus\{\tau\}.$ Additionally, since $\mathscr{I}^{\tau}_{i}(m)$ is independent of the ranking $m$, it remains constant for all $m$.
(c) $\forall i,j\in\mathcal{P}\setminus\{\tau\}$, since $d_{i}^{\tau}$ and $d_{j}^{\tau}$ are independent of $r_{ij}$, following the same proof techniques in (b), we have $\frac{\partial\text{PCS}\left(\tau\right)}{\partial r_{ij}}=2\cdot\frac{\partial\text{PCS}(\tau)}{\partial{{\widetilde{r}}_{i,j}}^{\tau}}\frac{\partial\widetilde{r}_{i,j}^{\tau}}{\partial r_{ij}}$. It is easy to verify that $\frac{\partial\widetilde{r}_{i,j}^{\tau}}{\partial r_{ij}}>0$ and according to (9), $\frac{\partial\text{PCS}(\tau)}{\partial{{\widetilde{r}}_{i,j}}^{\tau}}\geq 0$. Therefore, $\frac{\partial\text{PCS}\left(\tau\right)}{\partial r_{ij}}\geq 0$. Moreover, as for the order of magnitude, we have $\frac{\partial\text{PCS}\left(\tau\right)}{\partial r_{ij}}=\mathcal{O}(e^{-\mathscr{I}^{\tau}_{ij}N})$. If $\mathscr{I}^{\tau}_{i}>\mathscr{D}^{\tau}_{i}$ holds, then since $\mathscr{I}^{\tau}_{ij}\geq\mathscr{I}^{\tau}_{i}$, we have $|\frac{\partial\text{PCS}\left(\tau\right)}{\partial r_{ij}}|=o(\frac{\partial\text{PCS}\left(\tau\right)}{\partial r_{i\tau}})$ ($i,j\neq\tau$). The influence of $\frac{\partial\text{PCS}\left(\tau\right)}{\partial r_{ij}}$ is negligible compared to $\frac{\partial\text{PCS}\left(\tau\right)}{\partial r_{i\tau}}$.
Appendix B Proof of Theorem 3 and the Order of $\gamma$ in (3).
Figure 6: Starting from the initialization, Strategy $\mathcal{R}$ directly increases the total sample size, while Strategy $\mathcal{C}$ is equivalent to first increasing the total sample size and then changing the covariance.
{pf}
We view $\text{PCS}(\tau;\Sigma,N)$ as a function of the total sample size and covariance matrix. First, we consider the strategy $\mathcal{R}$: randomly assigning $p/k$ alternatives to each processor. In the following analysis, we focus on calculating the expected sample complexity reduction for the first processor. By symmetry, multiplying the result by $k$ gives the total sample complexity reduction. The index set of alternatives in the first processor is denoted as $\mathcal{P}_{1}$. The local best $\tau_{1}$} of this processor is assumed to be from cluster $\mathcal{G}_{1}$. In this processor, besides $\tau_{1}$, there are $\eta$ alternatives from cluster $\mathcal{G}_{1}$ (the index set of those alternatives is denoted as $\mathcal{J}_{1}\subseteq\mathcal{P}_{1}$). The remaining $p/k-\eta$ alternatives are from other clusters. $\eta$ follows a hypergeometric distribution $H(p,p/k,p/k)$ and $\mathbb{E}(\eta)=\frac{(p/k)^{2}}{p}$. The covariance matrix of alternatives in this processor is $\Sigma$. Suppose the initialization sample size is $N_{0}$ and the corresponding $\text{PCS}({\tau_{1}};\Sigma,N_{0})$ of this processor is $1-\alpha_{0}$ at this point. We continue sampling until PCS reaches $1-\alpha$. The sample size is increased to $N_{R}$. According to mean value theorem,
$\text{PCS}({\tau_{1}};\Sigma,N_{R})-\text{PCS}({\tau_{1}};\Sigma,N_{0})=\frac{\partial\text{PCS}\left({\tau_{1}};\Sigma,N_{0}+\xi_{R}(N_{R}-N_{0})\right)}{\partial N}\cdot(N_{R}-N_{0})=\alpha_{0}-\alpha,$ (19)
where $\xi_{R}\in(0,1)$. This process is illustrated in the red line in Figure 6.
Next we consider the strategy $\mathcal{C}$. By using correlation-based clustering, $\beta(p/k)$ alternatives in the first processor are correctly clustered (i.e., from $\mathcal{G}_{1}$), where $\beta=\frac{\Sigma_{i\in\mathcal{P}_{1}\setminus\{\tau_{1}\}}\mathbb{I}(G_{n}(i)=G(i))}{p/k}$ and $n$ is the sample size used for learning correlation information. Since $\mathrm{PCC}=P(\Pi_{n}=\Pi)\leq P(G_{n}(i)=G(i))$, $\forall i\in\mathcal{P}_{1}$, and the cardinality $|\mathcal{P}_{1}\setminus\{\tau_{1}\}|=p/k$, we have
$\mathbb{E}(\beta)=\mathbb{E}\bigg(\frac{\Sigma_{i\in\mathcal{P}_{1}\setminus\{\tau_{1}\}}\mathbb{I}(G_{n}(i)=G(i))}{p/k}\bigg)=\frac{\Sigma_{i\in\mathcal{P}_{1}\setminus\{\tau_{1}\}}P(G_{n}(i)=G(i))}{p/k}\geq\frac{p/k\cdot\mathrm{PCC}}{p/k}=\mathrm{PCC}.$ (20)
The transition from Strategy $\mathcal{R}$ to Strategy $\mathcal{C}$ is equivalent to increasing the number of alternatives belonging to cluster $\mathcal{G}_{1}$ from $\eta$ to $\beta(p/k)$. Furthermore, this is equivalent to changing the correlation structure of this processor from $\Sigma$ to $\Sigma^{\prime}$ while keeping the means and variances unchanged. Specifically, the change in the correlation structure is as follows: (i) The correlation coefficients between $\beta(p/k)-\eta$ alternatives in $\mathcal{P}_{1}\setminus\mathcal{J}_{1}$ and alternative $\tau_{1}$} are increased by $\Delta r$ (the index set of those $\beta(p/k)-\eta$ alternatives is denoted as $\mathcal{J}_{2}$). For simplifying calculations, we assume that among the remaining $\mathcal{J}_{3}\triangleq\mathcal{P}_{1}\setminus(\mathcal{J}_{1}\cup\mathcal{J}_{2}\cup\{\tau_{1}\})$, there are no alternatives belonging to the same cluster as $\mathcal{J}_{2}$. This is reasonable, because according to Lemma 1, the impact of the correlation between $\mathcal{J}_{2}$ and $\mathcal{J}_{3}$ is a negligible lower-order term. (ii) As for the correlation coefficients $\{r_{i,j}\}_{i,j\in\mathcal{P}_{1}\setminus\{{\tau_{1}}\}}$ between two alternatives in $\mathcal{P}_{1}\setminus\{{\tau_{1}}\}$, $\eta[\beta(p/k)-\eta]$ of them are increased by $\Delta r$ (i.e., the correlations between the alternatives in $\mathcal{J}_{1}$ and $\mathcal{J}_{2}$). The remaining correlation coefficients remain unchanged.
Starting from the initialization $N_{0}$ samples, we continue sampling until PCS reaches $1-\alpha$ and the total sample size is $N_{C}$. This process is depicted by the green line in Figure 6, where the correlation structure is first altered to $\Sigma^{\prime}$, and then the total sample size is increased to $N_{C}$. This process is equivalent to the one illustrated by the blue line in Figure 6, which involves first increasing total sample size, keeping the correlation coefficients constant and then altering the correlation structure. Then by mean value theorem,
$\text{PCS}({\tau_{1}};\Sigma^{\prime},N_{C})-\text{PCS}({\tau_{1}};\Sigma,N_{0})$ (21)
$=\frac{\partial\text{PCS}\left({\tau_{1}}\right)}{\partial N}\cdot(N_{C}-N_{0})+\sum_{i\in\mathcal{J}_{2}}\frac{\partial\text{PCS}\left({\tau_{1}}\right)}{\partial r_{i,{\tau_{1}}}}\Delta r+\sum_{i\in\mathcal{J}_{1},j\in\mathcal{J}_{2}}\frac{\partial\text{PCS}\left({\tau_{1}}\right)}{\partial r_{ij}}\Delta r=\alpha_{0}-\alpha,$
where $\frac{\partial\text{PCS}\left({\tau_{1}}\right)}{\partial N}$ is evaluated at a point $(\Sigma,N_{0}+\xi_{C}(N_{C}-N_{0}))$ and $\xi_{C}\in(0,1)$. $\frac{\partial\text{PCS}\left({\tau_{1}}\right)}{\partial r_{i{\tau_{1}}}}$ and $\frac{\partial\text{PCS}\left({\tau_{1}}\right)}{\partial r_{ij}}$ are evaluated at a point $(\Sigma+\xi^{\prime}(\Sigma^{\prime}-\Sigma),N_{0}+\xi_{C}(N_{C}-N_{0}))$, where $\xi^{\prime}\in(0,1)$. With Lemma 1, we have $\sum_{i\in\mathcal{J}_{1},j\in\mathcal{J}_{2}}\frac{\partial\text{PCS}\left({\tau_{1}}\right)}{\partial r_{ij}}\Delta r>0.$ Since each local best has maximum mean, according to Theorem 2 and Corollary 1, $\frac{\partial\text{PCS}\left({\tau_{1}}\right)}{\partial r_{i,{\tau_{1}}}}\geq 0$ for any $i\neq{\tau_{1}}$, if we omit lower-order terms. As shown in Figure 6, the process of strategy $\mathcal{R}$ and $\mathcal{C}$ have the same starting ($\alpha_{0}$) and ending points ($\alpha$). With 19 and 21, the identical $\alpha_{0}-\alpha$ terms are canceled out, then we have ∑i∈J2∂PCS(τ1)∂ri,τ1Δr≤∑i∈J2∂PCS(τ1)∂ri,τ1Δr+∑i∈J1, j∈J2∂PCS(τ1)∂rijΔr=∂PCS(τ1;Σ,N0+ξR(NR-N0))∂N⋅(NR-N0)-∂PCS(τ1;Σ,N0+ξC(NC-N0))∂N⋅(NC-N0)≤∂PCS(τ1;Σ,N0+ξC(NC-N0))∂N⋅(NR-NC).
The $\leq"$ in the last line arises because the PCS is a concave function that is increasing with respect to $N$, and $N_{R}\geq N_{C}$, thus $\frac{\partial\text{PCS}\left({\tau_{1}};\Sigma,N_{0}+\xi_{R}(N_{R}-N_{0})\right)}{\partial N}\leq\frac{\partial\text{PCS}\left({\tau_{1}};\Sigma,N_{0}+\xi_{C}(N_{C}-N_{0})\right)}{\partial N}$. Note that this is the only step where the concavity assumption is used. Without this assumption, the above inequality can be rewritten as $\leq\max\{\frac{\partial\text{PCS}\left({\tau_{1}};\Sigma,N_{0}+\xi_{C}(N_{C}-N_{0})\right)}{\partial N},\frac{\partial\text{PCS}\left({\tau_{1}};\Sigma,N_{0}+\xi_{R}(N_{R}-N_{0})\right)}{\partial N}\}\cdot(N_{R}-N_{C})$, and a similar final result can still be obtained.
[…]
Next, we prove that the condition $\gamma$ is at least $\mathcal{O}(1)$" is met for sampling strategies with lowest sample size growth rate $\mathcal{O}(p)$. According to Theorem 2 and Corollary 1, we have
$\frac{\partial\text{PCS}({\tau_{1}})}{\partial r_{i,{\tau_{1}}}}={\widetilde{D}}^{\tau}_{i}+o({\widetilde{D}}^{\tau}_{i}).$ (22)
Following the same proof techniques as Theorem 2 and taking the derivative of PCS with respect to individual sample size $N_{i}$, we have
$\frac{\partial\text{PCS}\left({\tau_{1}}\right)}{\partial N_{i}}=D^{\prime}_{i}+o(D^{\prime}_{i}),\ \ D^{\prime}_{i}=\frac{\partial d_{i}{}^{{\tau_{1}}}}{\partial N_{i}}\int_{-d_{1}^{\tau_{1}}}^{\infty}\cdots\int_{-d_{p}{}^{{\tau_{1}}}}^{\infty}f_{y^{{\tau_{1}}}}\left(y_{1},\cdots,-d_{i}{}^{{\tau_{1}}},\cdots,y_{p}\right)dy_{1}\cdots dy_{p}.$ (23)
Therefore, for any sampling strategy with lowest growth rate of total sample size $N=\mathcal{O}(p)$, we can easily conclude that $\frac{\partial\text{PCS}\left({\tau_{1}}\right)}{\partial N}$ is at most $\mathcal{O}(1)$. Then, by definition, we can easily conclude that $\gamma$ is at least $\mathcal{O}(1)$.
[…]
References in Appendices
  • Azaïs and Wschebor (2009) Azaïs, J.M., Wschebor, M., 2009. Level sets and extrema of random processes and fields. John Wiley & Sons.
  • Hashorva and Hüsler (2003) Hashorva, E., Hüsler, J., 2003. On multivariate gaussian tails. Annals of the Institute of Statistical Mathematics 55, 507–522.